### v2

# Distributional Reduction: Unifying Dimensionality Reduction and Clustering with Gromov-Wasserstein

Abstract
Unsupervised learning aims to capture the underlying structure of potentially large and high-dimensional datasets. Traditionally, this involves using dimensionality reduction (DR) methods to project data onto lower-dimensional spaces or organizing points into meaningful clusters (clustering). In this work, we revisit these approaches under the lens of optimal transport and exhibit relationships with the Gromov-Wasserstein problem. This unveils a new general framework, called distributional reduction, that recovers DR and clustering as special cases and allows addressing them jointly within a single optimization problem. We empirically demonstrate its relevance to the identification of low-dimensional prototypes representing data at different scales, across multiple image and genomic datasets.
1 Introduction
One major objective of unsupervised learning Hastie2009 is to provide interpretable and meaningful approximate representations of the data that best preserve its structure i.e. the underlying geometric relationships between the data samples. Similar in essence to Occam’s principle frequently employed in supervised learning, the preference for unsupervised data representation often aligns with the pursuit of simplicity, interpretability or visualizability in the associated model. These aspects are determinant in many real-world applications where the interaction with domain experts is paramount for interpreting the results and extracting meaningful insights from the model. For instance, the design of tissue atlases or the inference of single-cell trajectories, which are essential for addressing various biological challenges rao2021exploring ; saelens2019comparison , depends on the analysis of metacells, i.e. granular and interpretable representations of cells that enable separating sampling effect from biological variance baran2019metacell .
Dimensionality reduction and clustering. When faced with the question of extracting interpretable representations, from a dataset ${\mathbf{X}}=({\mathbf{x}}_{1},...,{\mathbf{x}}_{N})^{\top}\in\mathbb{R}^{N\times p}$ of $N$ samples in $\mathbb{R}^{p}$, the machine learning community has proposed a variety of methods. Among them, dimensionality reduction (DR) algorithms have been widely used to summarize data in a low-dimensional space ${\mathbf{Z}}=({\mathbf{z}}_{1},...,{\mathbf{z}}_{N})^{\top}\in\mathbb{R}^{N\times d}$ with $d\ll p$, allowing for visualization of every individual points for small enough $d$ lee2007nonlinear ; van2009dimensionality . Another major approach is to cluster the data into $n$ groups, with $n$ typically much smaller than $N$, and to summarize these groups through their centroids bishop2006pattern ; von2007tutorial . Clustering is particularly interpretable since it provides a smaller number of representative points that can be easily inspected. The cluster assignments can also be analyzed. Both DR and clustering follow a similar philosophy of summarization and reduction of the dataset using a smaller size representation.
Two sides of the same coin. As a matter of fact, methods from both families share many similitudes, including the construction of a similarity graph between input samples. In clustering, many popular approaches design a reduced or coarsened version of the initial similarity graph while preserving some of its spectral properties von2007tutorial ; schaeffer2007graph . In DR, the goal is to solve the inverse problem of finding low-dimensional embeddings that generate a similarity graph close to the one computed from input data points ham2004kernel ; hinton2002stochastic . Our work builds on these converging viewpoints and addresses the following question: can DR and clustering be expressed in a common and unified framework ?
Refer to caption Refer to caption
Figure 1: Illustration of our DistR method on two toys examples. Both pairs of representations ${\mathbf{X}},{\mathbf{Z}}$ are respectively associated to their similarity matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$, coupled via the matrix ${\mathbf{T}}$ (depicted in purple, with its marginals).
A distributional perspective. To answer this question, we propose to look at both problems from a distributional point of view, treating the data as an empirical probability distribution $\mu=\frac{1}{N}\sum_{i}\delta_{{\mathbf{x}}_{i}}$. This enables to consider statistical measures of similarity such as Optimal Transport (OT), which is at the core of our work. On the one hand, OT and clustering are strongly related. The celebrated k-means algorithm can be seen as a particular case of minimal Wasserstein estimator where a distribution of $n$ Diracs is optimized w.r.t their weights and positions Canas12 . Other connections between spectral clustering and the OT-based Gromov-Wasserstein (GW) distance have been recently developed in chowdhury2021generalized ; chen2023gromov ; vincent2021semi . On the other hand, the link between DR and OT has not been explored yet. DR methods, when modeling data as distributions, usually focus on joint distribution between samples within each space separately, see e.g. van2023snekhorn or lu2019doubly . Consequently, they do not consider couplings to transport samples across spaces of varying dimensions.
Contributions. In this paper, we bridge this gap by proposing a general distributional framework that encompasses both DR and clustering as special cases. We notably cast those problems as finding a reduced distribution that minimizes the GW divergence from the original empirical data distribution. Our method proceeds by first constructing an input similarity matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ that is matched with the embedding similarity ${\mathbf{C}}_{Z}({\mathbf{Z}})$ through an OT coupling matrix ${\mathbf{T}}$. The latter establishes correspondences between input and embedding samples. We illustrate this principle in Figure 1 where one can notice that ${\mathbf{C}}_{Z}({\mathbf{Z}})$ preserves the topology of ${\mathbf{C}}_{X}({\mathbf{X}})$ with a reduced number of nodes. The adaptivity of our model that can select an effective number of cluster $<n$, is visible in the bottom plot, where only the exact number of clusters in the original data ($9$ out of the $12$ initially proposed) is automatically recovered. Our method can operate in any embedding space, which is illustrated by projecting in either a 2D Euclidean plane or a Poincaré ball as embedding spaces.
Outline. We show that this framework is versatile and allows recovering many popular DR methods such as the kernel PCA and neighbor embedding algorithms, but also clustering algorithms such as weighted k-means and its kernel counterpart including spectral clustering chan2004optimization ; dhillon2007weighted . We first prove in Section 3 that DR can be formulated as a GW projection problem under some conditions on the loss and similarity functions. We then propose in Section 4 a novel formulation of data summarization as a minimal GW estimator that allows selecting both the dimensionality of the embedding $d$ (DR) and the cardinality of the support $n$ (Clustering). Finally, we show in section 5 the practical interest of our approach on images and genomics datasets.
Notations. The $i^{th}$ entry of a vector ${\mathbf{v}}$ is denoted as either $v_{i}$ or $[{\mathbf{v}}]_{i}$. Similarly, for a matrix ${\mathbf{M}}$, $M_{ij}$ and $[{\mathbf{M}}]_{ij}$ both denote its entry $(i,j)$. $S_{N}$ is the set of permutations of ${[\![N]\!]}$. $P_{N}(\mathbb{R}^{d})$ refers to the set of discrete probability measures composed of N points of $\mathbb{R}^{d}$. $\Sigma_{N}$ stands for the probability simplex of size $N$ that is $\Sigma_{N}:=\{{\mathbf{h}}\in\mathbb{R}^{N}_{+}\>\text{s.t.}\>\sum_{i}h_{i}=1\}$. $\log({\mathbf{M}})$ and $\exp({\mathbf{M}})$ are to be understood element-wise. For ${\mathbf{x}}\in\mathbb{R}^{N},\operatorname{diag}({\mathbf{x}})$ denotes the diagonal matrix whose elements are the $x_{i}$.
2 Background on Dimensionality Reduction and Optimal Transport
We first review the most popular DR approaches and introduce the Gromov-Wasserstein problem.
2.1 Unified View of Dimensionality Reduction
Let ${\mathbf{X}}=({\mathbf{x}}_{1},...,{\mathbf{x}}_{N})^{\top}\in\mathbb{R}^{N\times p}$ be an input dataset. Dimensionality reduction focuses on constructing a low-dimensional representation or embedding ${\mathbf{Z}}=({\mathbf{z}}_{1},...,{\mathbf{z}}_{N})^{\top}\in\mathbb{R}^{N\times d}$, where $d<p$. The latter should preserve a prescribed geometry for the dataset encoded via a symmetric pairwise similarity matrix ${\mathbf{C}}_{{\mathbf{X}}}\in\mathbb{R}_{+}^{N\times N}$ obtained from ${\mathbf{X}}$. To this end, most popular DR methods optimize ${\mathbf{Z}}$ such that a certain pairwise similarity matrix in the output space matches ${\mathbf{C}}_{{\mathbf{X}}}$ according to some criteria. We subsequently introduce the functions
${\mathbf{C}}_{X}:\mathbb{R}^{N\times p}\to\mathbb{R}^{N\times N},{\mathbf{C}}_{Z}:\mathbb{R}^{N\times d}\to\mathbb{R}^{N\times N}\,,$ (1)
which define pairwise similarity matrices in the input and output space, from the dataset ${\mathbf{X}}$ and the embedding ${\mathbf{Z}}$ respectively. The DR problem can be formulated quite generally as the optimization problem
$\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\>\sum_{(i,j)\in{[\![N]\!]}^{2}}L\big{(}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}\big{)}\,.$ (DR)
where $L:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ is a loss that quantifies the discrepancy between similarities between points in the input space $\mathbb{R}^{p}$ and in the output space $\mathbb{R}^{d}$. Various losses are used, such as the quadratic loss $L_{2}(x,y)\coloneqq(x-y)^{2}$ or the Kullback-Leibler divergence $L_{\operatorname{KL}}(x,y)\coloneqq x\log(x/y)-x+y$. Below, we recall several popular methods that can be placed within this framework.
Spectral methods. When ${\mathbf{C}}_{X}({\mathbf{X}})$ is a positive semi-definite matrix, eq. DR recovers spectral methods by choosing the quadratic loss $L=L_{2}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})=(\langle{\mathbf{z}}_{i},{\mathbf{z}}_{j}\rangle)_{(i,j)\in{[\![N]\!]}^{2}}$ the matrix of inner products in the embedding space. Indeed, in this case, the objective value of eq. DR reduces to
$\sum_{(i,j)\in{[\![N]\!]}^{2}}\!\!\!\!\!L_{2}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\langle{\mathbf{z}}_{i},{\mathbf{z}}_{j}\rangle)=\|{\mathbf{C}}_{X}({\mathbf{X}})-{\mathbf{Z}}{\mathbf{Z}}^{\top}\|^{2}_{F}\,$
where $\|\cdot\|_{F}$ is the Frobenius norm. This problem is commonly known as kernel Principal Component Analysis (PCA) scholkopf1997kernel and an optimal solution is given by ${\mathbf{Z}}^{\star}=(\sqrt{\lambda_{1}}{\mathbf{v}}_{1},...,\sqrt{\lambda_{d}}{\mathbf{v}}_{d})^{\top}$ where $\lambda_{i}$ is the $i$-th largest eigenvalue of ${\mathbf{C}}_{X}({\mathbf{X}})$ with corresponding eigenvector ${\mathbf{v}}_{i}$ eckart1936approximation . As shown by ham2004kernel ; ghojogh2021unified , numerous dimension reduction methods can be categorized in this manner. This includes PCA when ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ is the matrix of inner products in the input space; (classical) multidimensional scaling borg2005modern , when ${\mathbf{C}}_{X}({\mathbf{X}})=-\frac{1}{2}{\mathbf{H}}{\mathbf{D}}_{\mathbf{X}}{\mathbf{H}}$ with ${\mathbf{D}}_{\mathbf{X}}$ the matrix of squared euclidean distance between the points in $\mathbb{R}^{p}$ and ${\mathbf{H}}={\mathbf{I}}_{N}-\frac{1}{N}\bm{1}_{N}\bm{1}_{N}^{\top}$ is the centering matrix; Isomap tenenbaum2000global , with ${\mathbf{C}}_{X}({\mathbf{X}})=-\frac{1}{2}{\mathbf{H}}{\mathbf{D}}_{\mathbf{X}}^{(g)}{\mathbf{H}}$ with ${\mathbf{D}}_{\mathbf{X}}^{(g)}$ the geodesic distance matrix; Laplacian Eigenmap belkin2003laplacian , with ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{L}}_{\mathbf{X}}^{\dagger}$ the pseudo-inverse of the Laplacian associated to some adjacency matrix ${\mathbf{W}}_{\mathbf{X}}$; but also Locally Linear Embedding roweis2000nonlinear , and Diffusion Map coifman2006diffusion (for all of these examples we refer to (ghojogh2021unified, , Table 1)).
Neighbor embedding methods. An alternative group of methods relies on neighbor embedding techniques which consists in minimizing in ${\mathbf{Z}}$ the quantity
$\sum_{(i,j)\in{[\![N]\!]}^{2}}L_{\operatorname{KL}}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})\>.$ (NE)
Within our framework, this corresponds to eq. DR with $L=L_{\operatorname{KL}}$. The objective function of popular methods such as stochastic neighbor embedding (SNE) hinton2002stochastic or t-SNE van2008visualizing can be derived from eq. NE with a particular choice of ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$. For instance SNE and t-SNE both consider in the input space a symmetrized version of the entropic affinity vladymyrov2013entropic ; van2023snekhorn . In the embedding space, ${\mathbf{C}}_{Z}({\mathbf{Z}})$ is usually constructed from a “kernel” matrix ${\mathbf{K}}_{\mathbf{Z}}$ which undergoes a scalar van2008visualizing , row-stochastic hinton2002stochastic or doubly stochastic lu2019doubly ; van2023snekhorn normalization. Gaussian kernel $[{\mathbf{K}}_{\mathbf{Z}}]_{ij}=\exp(-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})$, or heavy-tailed Student-t kernel $[{\mathbf{K}}_{\mathbf{Z}}]_{ij}=(1+\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})^{-1}$, are typical choices van2008visualizing . We also emphasize that one can retrieve the UMAP objective mcinnes2018umap from eq. DR using the binary cross-entropy loss. A comprehensive overview and probabilistic analysis of these methods can be found in van2022probabilistic .
Remark 2.1.
The usual formulations of neighbor embedding methods rely on the loss $L(x,y)=x\log(x/y)$ instead of $L_{\operatorname{KL}}$. However, due to the normalization, the total mass $\sum_{ij}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}$ is constant (often equal to $1$) in all of the cases mentioned above. Thus the minimization in ${\mathbf{Z}}$ with the $L_{\operatorname{KL}}$ formulation is equivalent to the usual formulations of neighbor embedding objectives.
Non-Euclidean geometries. Most DR methods can also be extended to incorporate non-Euclidean geometries. For instance, Hyperbolic spaces Chami21 ; Fan_2022_CVPR ; Guo22 ; Lin23 are of particular interest as they can capture hierarchical structures more effectively than Euclidean spaces and further mitigate the curse of dimensionality. However, these methods introduce additional hyperparameters and optimization difficulties that can limit their applicability. For this reason, we report some numerical experiments with such kernels in Section E.7 as proofs of concept supporting the versatility of our methods.
2.2 Optimal Transport Across Spaces
Optimal Transport (OT) villani2009optimal ; peyre2019computational is a popular framework for comparing probability distributions and is at the core of our contributions. We review in this section the Gromov-Wasserstein formulation of OT aiming at comparing distributions across spaces.
Gromov-Wasserstein (GW). The GW framework memoli2011gromov ; sturm2012space comprises a collection of OT methods designed to compare distributions by examining the pairwise relations within each space. For two matrices ${\mathbf{C}}\in\mathbb{R}^{N\times N}$, $\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}$, and weights ${\mathbf{h}}\in\Sigma_{N},\overline{{\mathbf{h}}}\in\Sigma_{n}$, the GW discrepancy with inner loss $L$ peyre2016gromov is defined as
$\begin{split}&\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=\min_{{\mathbf{T}}\in{\mathcal{U}}({\mathbf{h}},\overline{{\mathbf{h}}})}E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})\ =\min_{{\mathbf{T}}\in{\mathcal{U}}({\mathbf{h}},\overline{{\mathbf{h}}})}\sum_{ijkl}L(C_{ij},\overline{C}_{kl})T_{ik}T_{jl}\,,\end{split}$ (GW)
and ${\mathcal{U}}({\mathbf{h}},\overline{{\mathbf{h}}})=\left\{{\mathbf{T}}\in\mathbb{R}_{+}^{N\times n}:{\mathbf{T}}\bm{1}_{n}={\mathbf{h}},{\mathbf{T}}^{\top}\bm{1}_{N}=\overline{{\mathbf{h}}}\right\}$ is the set of couplings between ${\mathbf{h}}$ and $\overline{{\mathbf{h}}}$. In this formulation, both pairs $({\mathbf{C}},{\mathbf{h}})$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ can be interpreted as graphs with corresponding connectivity matrices ${\mathbf{C}},\overline{\mathbf{C}}$, and where nodes are weighted by histograms ${\mathbf{h}},\overline{\mathbf{h}}$ (with implicit supports). Equation GW is thus a quadratic problem (in ${\mathbf{T}}$) which consists in finding a soft-assignment matrix ${\mathbf{T}}$ that aligns the nodes of the two graphs in a way that preserves their pairwise connectivities.
From a distributional perspective, GW can also be viewed as a distance between distributions that do not belong to the same metric space. For two discrete probability distributions $\mu_{X}=\sum_{i=1}^{N}[{\mathbf{h}}_{X}]_{i}\delta_{{\mathbf{x}}_{i}}\in{\mathcal{P}}_{N}(\mathbb{R}^{p}),\mu_{Z}=\sum_{i=1}^{n}[{\mathbf{h}}_{Z}]_{i}\delta_{{\mathbf{z}}_{i}}\in{\mathcal{P}}_{n}(\mathbb{R}^{d})$ and pairwise similarity matrices ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$ associated with the explicit supports ${\mathbf{X}}=({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{n})^{\top}$ and ${\mathbf{Z}}=({\mathbf{z}}_{1},\cdots,{\mathbf{z}}_{n})^{\top}$, the quantity $\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{h}}_{X},{\mathbf{h}}_{Z})$ is a measure of dissimilarity or discrepancy between $\mu_{X},\mu_{Z}$. Specifically, when $L=L_{2}$, and ${\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}})$ are pairwise distance matrices, GW defines a proper distance between $\mu_{X}$ and $\mu_{Z}$ with respect to measure preserving isometries [^1] With weaker assumptions on ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$, GW defines a pseudo-metric w.r.t. a different notion of isomorphism chowdhury2019gromov . See Appendix B for more details. memoli2011gromov .
Due to its versatile properties, notably in comparing distributions over different domains, the GW problem has found many applications in machine learning, e.g., for 3D meshes alignment solomon2016entropic ; ezuz2017gwcnn , NLP alvarez2018gromov , (co-)clustering peyre2016gromov ; redko2020co , single-cell analysis demetci2020gromov , neuroimaging thual2022aligning , graph representation learning xu2020gromov ; vincent2021online ; liu2022robust ; vincent2022template ; pmlr-v202-zeng23c and partitioning xu2019scalable ; chowdhury2021generalized . In this work, we leverage the GW discrepancy to extend classical DR approaches, framing them as the projection of a distribution onto a space of lower dimensionality.
3 Optimal Transport Formulation of DR
In this section, we outline the strong connections between the classical eq. DR and the GW problem.
Gromov-Monge interpretation of DR. As suggested by eq. DR, dimension reduction seeks to find embeddings ${\mathbf{Z}}$ so that the similarity between the $(i,j)$ samples of the input data is as close as possible to the similarity between the $(i,j)$ samples of the embeddings. Under reasonable assumptions about ${\mathbf{C}}_{Z}$, this also amounts to identifying the embedding ${\mathbf{Z}}$ and the best permutation that realigns the two similarity matrices ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$. Recall that the function ${\mathbf{C}}_{Z}$ is equivariant by permutation, if, for any $N\times N$ permutation matrix ${\mathbf{P}}$ and any ${\mathbf{Z}}$, ${\mathbf{C}}_{Z}({\mathbf{P}}{\mathbf{Z}})={\mathbf{P}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{P}}^{\top}$ bronstein2021geometric . This type of assumption is natural for ${\mathbf{C}}_{Z}$: if we rearrange the order of samples (i.e. the rows of ${\mathbf{Z}}$), we expect the similarity matrix between the samples to undergo the same rearrangement.
Lemma 3.1.
Let ${\mathbf{C}}_{Z}$ be a permutation equivariant function and $L$ any loss. The minimum of eq. DR is equal to
$\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\min_{\sigma\in S_{N}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{\sigma(i)\sigma(j)})\>.$ (2)
Also, any solution ${\mathbf{Z}}$ of eq. DR is such that $({\mathbf{Z}},\operatorname{id})$ is solution of eq. 2. And conversely any $({\mathbf{Z}},\sigma)$ solution of eq. 2 is such that ${\mathbf{Z}}$ is a solution of eq. DR up to $\sigma$.
Lemma 3.1, proven in Section A.1, establishes an equivalence between eq. DR and the optimization of the embedding ${\mathbf{Z}}$ w.r.t a Gromov-Monge discrepancy [^2] Precisely $\min_{\sigma\in S_{N}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{\sigma(i)\sigma(j)})$ is the Gromov-Monge discrepancy between two discrete distributions, with the same number of atoms and uniform weights. memoli2018gromov given in eq. 2, which seeks to identify the permutation $\sigma$ that best aligns two similarity matrices, by solving a combinatorial quadratic assignment problem cela2013quadratic . We can delve deeper into these comparisons and demonstrate that the general formulation of dimension reduction is also equivalent to minimizing the Gromov-Wasserstein objective, which serves as a relaxation of the Gromov-Monge problem memoli2022comparison .
DR as GW Minimization. We suppose that the distributions have the same number of points ($N=n$) and uniform weights (${\mathbf{h}}_{Z}={\mathbf{h}}_{X}=\frac{1}{N}\bm{1}_{N}$). We recall that a matrix ${\mathbf{C}}\in\mathbb{R}^{N\times N}$ is conditionally positive definite (CPD), resp. conditionally negative definite (CND), if it is symmetric and $\forall{\mathbf{x}}\in\mathbb{R}^{N},{\mathbf{x}}^{\top}\mathbf{1}_{N}=0\text{s.t. }{\mathbf{x}}^{\top}{\mathbf{C}}{\mathbf{x}}\geq 0$, resp. $\leq 0$.
We thus extend Lemma 3.1 to the GW problem in the following (proof in Section A.2):
Theorem 3.2.
The minimum of eq. DR is equal to $\min_{{\mathbf{Z}}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ if:
  1. (i)
    (spectral methods) ${\mathbf{C}}_{X}({\mathbf{X}})$ is any matrix, $L=L_{2}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$.
  2. (ii)
    (neighbor embedding methods) $\operatorname{Im}({\mathbf{C}}_{X})\subseteq\mathbb{R}_{>0}^{N\times N}$, $L=L_{\operatorname{KL}}$, the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD and, for any ${\mathbf{Z}}$,
    ${\mathbf{C}}_{Z}({\mathbf{Z}})=\operatorname{diag}(\bm{\alpha}_{\mathbf{Z}}){\mathbf{K}}_{\mathbf{Z}}\operatorname{diag}(\bm{\beta}_{\mathbf{Z}})\,,$ (3)
    where $\bm{\alpha}_{\mathbf{Z}},\bm{\beta}_{\mathbf{Z}}\in\mathbb{R}^{N}_{>0}$ and ${\mathbf{K}}_{\mathbf{Z}}\in\mathbb{R}^{N\times N}_{>0}$ is such that $\log({\mathbf{K}}_{\mathbf{Z}})$ is CPD.
Remarkably, the first item of Theorem 3.2 shows that all spectral DR methods can be seen as OT problems in disguise, as they all equivalently minimize a GW problem. The second item also provides some insights into this equivalence in the case of neighbor embedding methods that require more assumptions. For instance, the Gaussian kernel ${\mathbf{K}}_{\mathbf{Z}}$, used extensively in DR (Section 2.1), satisfies the hypothesis as $\log({\mathbf{K}}_{\mathbf{Z}})=(-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})_{ij}$ is CPD (see e.g. maron2018probably ). The terms $\bm{\alpha}_{\mathbf{Z}},\bm{\beta}_{\mathbf{Z}}$ also allow for considering all the usual normalizations of ${\mathbf{K}}_{\mathbf{Z}}$: by a scalar so as to have $\sum_{ij}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}=1$, but also any row or doubly stochastic normalizations (with the Sinkhorn-Knopp algorithm sinkhorn1967concerning ).
Matrices satisfying $\log({\mathbf{K}}_{\mathbf{Z}})$ being CPD are well-studied in the literature and are known as infinitely divisible matrices bhatia2006infinitely . It is noteworthy that the t-Student kernel does not fall into this category. Moreover, in the aforementioned neighbor embedding methods, the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is generally not CPD. The intriguing question of generalizing this result with weaker assumptions on ${\mathbf{C}}_{Z}$ and ${\mathbf{C}}_{X}$ remains open for future research. Interestingly, we have observed in the numerical experiments performed in Section 5 that the symmetric entropic affinity of van2023snekhorn was systematically CPD.
Remark 3.3.
In Corollary A.3 of the appendix we also provide other sufficient conditions for neighbor embedding methods with the cross-entropy loss $L(x,y)=x\log(x/y)$. They rely on specific structures for ${\mathbf{C}}_{Z}$ but do not impose any assumptions on ${\mathbf{C}}_{X}$. Additionally, in Section A.3, we provide a necessary condition based on a bilinear relaxation of the GW problem. Although its applicability is limited due to challenges in proving it in full generality, it requires minimal assumptions on ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$ and $L$.
In essence, both Lemma 3.1 and Theorem 3.2 indicate that dimensionality reduction can be reframed from a distributional perspective, with the search for an empirical distribution that aligns with the data distribution in the sense of optimal transport, through the GW framework.In other words, DR is informally solving $\min_{{\mathbf{z}}_{1},\cdots,{\mathbf{z}}_{N}}\operatorname{GW}(\frac{1}{n}\sum_{i=1}^{N}\delta_{{\mathbf{x}}_{i}},\frac{1}{n}\sum_{i=1}^{N}\delta_{{\mathbf{z}}_{i}})$.
4 Distributional Reduction Framework
The above perspective on DR allows for the two following generalizations. Firstly, beyond solely determining the positions ${\mathbf{z}}_{i}$ of Diracs (as in classical DR) we can now optimize the mass of the distribution $\mu_{Z}$. This is interpreted as finding the relative importance of each point in the embedding ${\mathbf{Z}}$. More importantly, due to the flexibility of GW, we can also seek a distribution in the embedding with a smaller number of points $n<N$. This will result in both reducing the dimension and clustering the points in the embedding space through the optimal coupling. Informally, we thus propose a new Distributional Reduction (DistR) framework that aims at solving $\min_{\mu_{Z}\in{\mathcal{P}}_{n}(\mathbb{R}^{d})}\operatorname{GW}(\frac{1}{n}\sum_{i=1}^{N}\delta_{{\mathbf{x}}_{i}},\mu_{Z})$.
4.1 Distributional Reduction Problem
Precisely, the novel optimization problem that we tackle in this paper can be formulated as follows
$\min_{{\mathbf{Z}}\in\mathbb{R}^{n\times d}}\min_{{\mathbf{h}}_{Z}\in\Sigma_{n}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{h}}_{X},{\mathbf{h}}_{Z})$ (DistR)
This problem comes down to learning the graph $({\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{h}}_{Z})$ parametrized by ${\mathbf{Z}}$ that is the closest from $({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{h}}_{X})$ in the GW sense. When $n<N$, the embeddings then act as low-dimensional prototypical examples of input samples, whose learned relative importance ${\mathbf{h}}_{Z}$ accommodates clusters of varying proportions in the input data ${\mathbf{X}}$ (see Figure 1). We refer to them as prototypes. The weight vector ${\mathbf{h}}_{X}$ is typically assumed to be uniform, that is ${\mathbf{h}}_{X}=\frac{1}{N}\bm{1}_{N}$, in the absence of prior knowledge. As discussed in Section 3, traditional DR amounts to setting $n=N,{\mathbf{h}}_{Z}=\frac{1}{N}\bm{1}_{N}$.
One notable aspect of our model is its capability to simultaneously perform DR and clustering. Indeed, the optimal coupling ${\mathbf{T}}\in[0,1]^{N\times n}$ of problem eq. DistR is, by construction, a soft-assignment matrix from the input data to the embeddings. It allows each point ${\mathbf{x}}_{i}$ to be linked to one or more prototypes ${\mathbf{z}}_{j}$ (clusters). In Section 4.2 we explore conditions where these soft assignments transform into hard ones, such that each point is therefore linked to a unique prototype/cluster.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: GW projections of a single-cell genomics dataset chen2019high on regular grids with increasing resolutions, respectively encoded as $C_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ and $C_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$. Pixels on cropped grids are colored by interpolating samples’ colors according to the transport plan and their intensity is proportional to their mass.
A semi-relaxed objective.
For a given embedding ${\mathbf{Z}}$ and $L=L_{2}$, it is known that minimizing the DistR objective w.r.t ${\mathbf{h}}_{Z}$ is equivalent to a problem that is computationally simpler than the usual GW one, namely the semi-relaxed GW divergence $\operatorname{srGW}_{L}$ vincent2021semi :
$\min_{{\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{T}})\>,$ (srGW)
where ${\mathcal{U}}_{n}({\mathbf{h}}_{X}):=\left\{{\mathbf{T}}\in\mathbb{R}_{+}^{N\times n}:{\mathbf{T}}\bm{1}_{n}={\mathbf{h}}_{X}\right\}$. To efficiently address eq. DistR, we first observe that this equivalence holds for any inner divergence $L$ with a straightforward adaptation of the proof in vincent2021semi . Additionally, we prove that $\operatorname{srGW}_{L}$ remains a divergence as soon as $L$ is itself a divergence. Consequently, $\operatorname{srGW}_{L}$ vanishes iff both compared measures are isomorphic in a weak sense chowdhury2019gromov . We emphasize that taking a proper divergence $L$ is important (and basic assumptions on ${\mathbf{X}}$), as it avoids some trivial solutions as detailed in Appendix B.
Interestingly, srGW projections, i.e. optimizing only the weights ${\mathbf{h}}_{Z}$ over simple fixed supports ${\mathbf{Z}}$, have already remarkable representational capability. We illustrate this in Figure 2, by considering projections of a real-world dataset over 2D grids of increasing resolutions. Setting ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ and ${\mathbf{C}}_{Z}({\mathbf{X}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$, we can see that those projections recover faithful coarsened representations of the embeddings learned using PCA. DistR aims to exploit the full potential of this divergence by learning a few optimal prototypes that best represent the dataset.
Computation. DistR is a non-convex problem that we propose to tackle using a Block Coordinate Descent algorithm (BCD, tseng2001convergence ) guaranteed to converge to local optimum grippo2000convergence ; Lyu2023bmm . The BCD alternates between the two following steps. First, we optimize in ${\mathbf{Z}}$ for a fixed transport plan using gradient descent with adaptive learning rates kingma2014adam . Then we solve for a srGW problem given ${\mathbf{Z}}$. To this end, we extended both solvers proposed in vincent2021semi to support losses $L_{2}$ and $L_{\operatorname{KL}}$. Namely, the Mirror Descent promoting smoothness along iterations depending on a hyparameter $\varepsilon>0$ and the Conditional Gradient seen as the edge-case $\varepsilon=0$. Following Proposition 1 in peyre2016gromov , a vanilla implementation leads to $\mathcal{O}(nN^{2}+n^{2}N)$ operations to compute the loss or its gradient. However in many DR methods, ${\mathbf{C}}_{X}({\mathbf{X}})$ or ${\mathbf{C}}_{Z}({\mathbf{Z}})$, or their transformations within the loss $L$, admit explicit low-rank factorizations. Including e.g. matrices involved in spectral methods and other similarity matrices derived from squared Euclidean distance matrices scetbon2022linear . In these settings, we exploit these factorizations to reduce the computational complexity of our solvers down to $\mathcal{O}(Nn(p+d)+(N+n)pd+min(n^{2},nd^{2}))$ when $L=L_{2}$, and $O(Nnd+n^{2}d)$ when $L=L_{\operatorname{KL}}$. We refer the reader interested in these algorithmic details to Appendix D. As shown in Section E.6, DistR runs in a few seconds on datasets with several thousands of samples.
Related work. The CO-Optimal Transport (COOT) clustering approach proposed in redko2020co is the closest to our work. As detailed in Section E.2, it simultaneously estimates sample and feature clustering and thus can be directly applied to perform joint clustering and DR. Despite SOTA performances on co-clustering tasks, COOT-clustering corresponds to a linear DR method which is limiting for joint clustering and DR tasks. In Section E.2, we show that COOT leads to a low average homogeneity of the prototypes $\{{\mathbf{z}}_{k}\}_{k\in{[\![n]\!]}}$, indicating that it often assigns points with different labels to the same prototype. In contrast, DistR exploits the more expressive non-linear similarity functions offered by existing DR methods and leads to significantly improved results. Other joint DR-clustering approaches, such as liu2022joint , involve modeling latent variables by a mixture of distributions. In comparison, DistR is more versatile, as it easily adapts to any $(L,{\mathbf{C}}_{X},{\mathbf{C}}_{Z})$ of existing DR methods.
4.2 Clustering Properties
We elaborate now on the links between DistR and clustering methods. In what follows, we call a coupling ${\mathbf{T}}\in[0,1]^{N\times n}$ with a single non-null element per row a membership matrix. When the coupling is a membership matrix each data point is associated with a single prototype thus achieving a hard clustering of the input samples. We will see that a link can be drawn with kernel k-means using the analogy of GW barycenters. More precisely the srGW barycenter vincent2021semi seeks for a closest target graph $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ from $({\mathbf{C}}_{X},{\mathbf{h}}_{X})$ by solving
$\min_{\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}}\min_{{\mathbf{T}}\in\mathcal{U}_{n}({\mathbf{h}}_{X})}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})\>.$ (srGWB)
We stress that the only (important) difference between eq. srGWB and eq. DistR is that there is no constraint imposed on $\overline{{\mathbf{C}}}$ in srGWB. In contrast, eq. DistR looks for minimizing over $\overline{{\mathbf{C}}}\in\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$. For instance, choosing ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$ in eq. DistR is equivalent to enforcing $\operatorname{rank}(\overline{{\mathbf{C}}})\leq d$ in eq. srGWB.
We establish below that srGWB is of particular interest for clustering. The motivation for this arises from the findings of chen2023gromov , which demonstrate that, when ${\mathbf{C}}_{X}({\mathbf{X}})$ is positive semi-definite and ${\mathbf{T}}$ is constrained to belong to the set of membership matrices (as opposed to couplings in ${\mathcal{U}}_{n}({\mathbf{h}})$), eq. srGWB is equivalent to a kernel k-means whose samples are weighted by ${\mathbf{h}}_{X}$ dhillon2004kernel ; dhillon2007weighted . Interestingly, these additional constraints are unnecessary. Indeed, we show below that the original srGWB problem admits membership matrices as the optimal coupling for a broader class of ${\mathbf{C}}_{X}({\mathbf{X}})$ input matrices (see proof in Appendix C).
Theorem 4.1.
Let ${\mathbf{h}}_{X}\in\Sigma_{N}$ and $L=L_{2}$. Suppose that for any ${\mathbf{X}}\in\mathbb{R}^{N\times p}$ the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD or CND. Then the problem eq. srGWB admits a membership matrix as optimal coupling.
Theorem 4.1 and relations proven in chen2023gromov state that eq. srGWB is equivalent to the aforementioned kernel k-means when ${\mathbf{C}}_{X}({\mathbf{X}})$ is positive semi-definite. Moreover, as the (hard) clustering property holds for more generic types of matrices, namely CPD and CND, srGWB stands out as a fully-fledged clustering method. Although these results do not apply directly to DistR, except if e.g. the dimension $d$ is set to the unknown rank of the srGW barycenter, we argue that they further legitimize the use of GW projections for clustering. Interestingly, we also observe in practice that the couplings obtained by DistR are always membership matrices, regardless of ${\mathbf{C}}_{Z}$. Further research need to be carried out to better understand this observation.
5 Numerical Experiments
In this section, we demonstrate the relevance of our approach for joint clustering and DR, over 8 labeled datasets detailed in Appendix E including: 3 image datasets (COIL-20 nene1996columbia , MNIST & fashion-MNIST xiao2017fashion ) and 5 genomic ones (PBMC wolf2018scanpy , SNA 1 & 2 chen2019high and ZEISEL 1 & 2zeisel2015cell ). To this end, we examine how effectively DistR prototypes offer discriminative representations of input samples at various granularities, using an evaluation system described below.
[…]
Evaluation metrics. We consider several evaluation metrics to quantify whether the output prototypes provide a meaningful summarized representation of the input datasets. As often done in the literature, we focus on supervised evaluations while assessing two main properties. Firstly, we wonder whether the prototypes are individually discriminating input samples $\{{\mathbf{x}}_{i}\}$ w.r.t their class labels $\{y_{i}\}$ via their associated prototype. We thus compute the homogeneity score rosenberg2007v that quantifies to which extent each prototype ${\mathbf{z}}_{k}$ is associated with points in the same class. Secondly, we aim to evaluate whether the embedding space is both globally discriminant, i.e. that prototypes organize themselves to form clusters that match the classes, measured by two metrics. i) We associate with each prototype ${\mathbf{z}}_{k}$ a label using majority voting from the associated samples. Then we make use of the popular silhouette score rousseeuw1987silhouettes to evaluate whether prototypes’ positions are correctly aligned with these labels. ii) As commonly done to assess DR representations huang2022towards , we also consider the downstream task of clustering the prototypes using the k-means algorithm. We then assign to each $\{{\mathbf{x}}_{i}\}$ the cluster of its associated prototype, acting as a predicted label, and compute the Normalized Mutual Information (NMI) kvaalseth2017normalized between these labels and $\{y_{i}\}$. More details about evaluation metrics can be found in Appendix E.
Refer to caption Refer to caption
Figure 3: Best trade-off between homogeneity vs silhouette (2 first plots), and homogeneity vs NMI (2 last plots). Scores are normalized in $\left[0,1\right]$ via min-max scaling over a dataset. Small markers represent scores for 5 runs for a given dataset, while big ones are their mean. For each method we illustrate the 20-80% percentiles of normalized scores as a colored surface.
Figure 4: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{2}$) produced by DistR using the t-SNE model: SEA similarity for ${\mathbf{C}}_{X}$ van2023snekhorn , Student’s kernel for ${\mathbf{C}}_{Z}$ and loss $L_{\operatorname{KL}}$.
Figure 5: Examples of 2-dimensional embeddings produced by DistR using the SEA similarity for ${\mathbf{C}}_{X}$ and the Student’s kernel for ${\mathbf{C}}_{Z}$. The latter computed in $\mathbb{R}^{2}$ for the first three datasets and the Poincaré ball for the last one. Displayed images are medoids for each cluster i.e. $\operatorname*{arg\,max}_{i}[{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}_{:,k}]_{i}$ for cluster $k$. The area of image $k$ is proportional to $[{\mathbf{h}}_{Z}]_{k}$.
[…]
6 Conclusion
By making a connection between the GW problem and popular clustering and DR algorithms, we proposed a unifying framework denoted as DistR. DistR enables transforming any DR algorithm into a joint DR-clustering method that produces embeddings (or prototypes) with chosen granularity. We believe that the versatility of the GW framework will enable new extensions in unsupervised learning. For instance, the formalism associated with (semi-relaxed) GW barycenters naturally enables considering multiple inputs, potentially unaligned and of different sizes. Hence our approach can be useful to address challenges related to the multi-view dimensionality reduction and clustering problems. Other promising directions involve, better capturing multiple dependency scales in the input data by hierarchically adapting the resolution of the embedding similarity graph, or enabling batch optimization of embeddings to operate over much larger datasets.
Appendix A Proof of results in Section 3
A.1 Proof of lemma 3.1
We recall the result. See 3.1
Proof.
By suboptimality of $\sigma=\operatorname{id}$ we clearly have
$\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\min_{\sigma\in S_{N}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{\sigma(i)\sigma(j)})\leq\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}).$ (4)
For the other direction, take an optimal solution $({\mathbf{Z}},\sigma)$ of eq. 2. Using the permutation equivariance of ${\mathbf{C}}_{Z}$, $[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{\sigma(i)\sigma(j)}=[{\mathbf{P}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{P}}^{\top}]_{ij}=[{\mathbf{C}}_{Z}({\mathbf{P}}{\mathbf{Z}})]_{ij}$ for some permutation matrix ${\mathbf{P}}$. But ${\mathbf{P}}{\mathbf{Z}}$ is admissible for problem eq. DR. Hence
$\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\min_{\sigma\in S_{N}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{\sigma(i)\sigma(j)})\geq\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}).$ (5)
A.2 Proof of theorem 3.2
In the following $\operatorname{DS}$ is the space of $N\times N$ doubly stochastic matrices. We begin by proving the first point of theorem 3.2. We will rely on the simple, but useful, result below
Proposition A.1.
Let $\Omega\subseteq\mathbb{R}$ and $\operatorname{Im}({\mathbf{C}}_{X})\subseteq\Omega^{N\times N}$. Suppose that $L(a,\cdot)$ is convex for any $a\in\Omega$ and
$\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})\leq\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d},{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})\,.$ (6)
Then the minimum eq. DR is equal to $\min_{{\mathbf{Z}}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$.
Proof.
Consider any doubly stochastic matrix ${\mathbf{T}}$ and note that $[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij}=\sum_{kl}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl}T_{ik}T_{jl}$. Using the convexity of $L(a,\cdot)$ for any $a\in\Omega$ and Jensen’s inequality we have
$\begin{split}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})&=\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\sum_{kl}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl}T_{ik}T_{jl})\\ &\leq\sum_{ijkl}L([{\mathbf{C}}_{\mathbf{X}}]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.\end{split}$ (7)
In particular
$\min_{{\mathbf{Z}}}\min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})\leq\min_{{\mathbf{Z}}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\sum_{ijkl}L([{\mathbf{C}}_{\mathbf{X}}]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.$ (8)
Hence, using eq. 6,
$\min_{{\mathbf{Z}}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})\leq\min_{{\mathbf{Z}}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\sum_{ijkl}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.$ (9)
But the converse inequality is also true by sub-optimality of ${\mathbf{T}}={\mathbf{I}}_{N}$ for the problem $\min_{{\mathbf{Z}}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\sum_{ijkl}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}$. Overall
$\min_{{\mathbf{Z}}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})=\min_{{\mathbf{Z}}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\sum_{ijkl}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.$ (10)
Now we conclude by using that the RHS of this equation is equivalent to the minimization in ${\mathbf{Z}}$ of $\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ (both problems only differ from a constant scaling factor $N^{2}$). ∎
As a consequence we have the following result.
Proposition A.2.
Let $\Omega\subseteq\mathbb{R}$ and $\operatorname{Im}({\mathbf{C}}_{X})\subseteq\Omega^{N\times N}$. The minimum eq. DR is equal to $\min_{{\mathbf{Z}}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$ when:
  1. (i)
    $L(a,\cdot)$ is convex for any $a\in\Omega$ and the image of ${\mathbf{C}}_{Z}$ is stable by doubly-stochastic matrices, i.e. ,
    $\forall{\mathbf{Z}}\in\mathbb{R}^{N\times d},\forall{\mathbf{T}}\in\operatorname{DS},\exists{\mathbf{Y}}\in\mathbb{R}^{N\times d},\ {\mathbf{C}}_{Z}({\mathbf{Y}})={\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}\,.$ (11)
  2. (ii)
    $L(a,\cdot)$ is convex and non-decreasing for any $a\in\Omega$ and
    $\forall{\mathbf{Z}}\in\mathbb{R}^{N\times d},\forall{\mathbf{T}}\in\operatorname{DS},\exists{\mathbf{Y}}\in\mathbb{R}^{N\times d},\ {\mathbf{C}}_{Z}({\mathbf{Y}})\leq{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}\,,$ (12)
    where $\leq$ is understood element-wise, i.e. , $\bm{A}\leq\bm{B}\iff\forall(i,j),\ A_{ij}\leq B_{ij}$.
Proof.
For the first point it suffices to see that the condition eq. 11 implies that $\{{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}:{\mathbf{Z}}\in\mathbb{R}^{N\times d},{\mathbf{T}}\in\operatorname{DS}\}\subseteq\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$ (in fact we have equality by choosing ${\mathbf{T}}=\mathbf{I}_{N}$) and thus eq. 6 holds and we apply proposition A.1.
For the second point we will also show that eq. 6 holds. Consider ${\mathbf{Z}}^{\star},{\mathbf{T}}^{\star}$ minimizers of $\min_{{\mathbf{Z}},{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})$. By hypothesis there exists ${\mathbf{Y}}\in\mathbb{R}^{N\times d}$ such that $\forall(i,j)\in{[\![N]\!]}^{2},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij}\geq[{\mathbf{C}}_{Z}({\mathbf{Y}})]_{ij}$. Since $L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\cdot)$ is non-decreasing for any $(i,j)$ then $\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij})\geq\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Y}})]_{ij})$ and thus $\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij})\geq\min_{{\mathbf{Z}}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})$ which gives the condition eq. 6 and we have the conclusion by proposition A.1. ∎
We recall that a function $R:\mathbb{R}^{N\times d}\to\mathbb{R}$ is called permutation invariant if $R({\mathbf{P}}{\mathbf{Z}})=R({\mathbf{Z}})$ for any ${\mathbf{Z}}\in\mathbb{R}^{N\times d}$ and $N\times N$ permutation matrix ${\mathbf{P}}$. From the previous results we have the following corollary, which proves, in particular, the first point of theorem 3.2.
Corollary A.3.
We have the following equivalences:
  1. (i)
    Consider the spectral methods which correspond to ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{Z}}{\mathbf{Z}}^{\top}$ and $L=L_{2}$. Then for any ${\mathbf{C}}_{X}$
    $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}L_{2}([{\mathbf{C}}_{\mathbf{X}}]_{ij},\langle{\mathbf{z}}_{i},{\mathbf{z}}_{j}\rangle)$ (13)
    and
    $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{L_{2}}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{Z}}{\mathbf{Z}}^{\top},\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$ (14)
    are equal.
  2. (ii)
    Consider the cross-entropy loss $L(x,y)=x\log(x/y)$ and ${\mathbf{C}}_{X}$ such that $\operatorname{Im}({\mathbf{C}}_{X})\subseteq\mathbb{R}_{+}^{N\times N}$. Suppose that the similarity in the output space can be written as
    $\forall(i,j)\in{[\![N]\!]}^{2},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}=f({\mathbf{z}}_{i}-{\mathbf{z}}_{j})/R({\mathbf{Z}})\,,$ (15)
    for some logarithmically concave function $f:\mathbb{R}^{d}\to\mathbb{R}_{+}$ and normalizing factor $R:\mathbb{R}^{N\times d}\to\mathbb{R}_{+}^{*}$ which is both convex and permutation invariant. Then,
    $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})$ (16)
    and
    $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$ (17)
    are equal.
Proof.
For the first point we show that the condition eq. 11 of proposition A.2 is satisfied. Indeed take any ${\mathbf{Z}},{\mathbf{T}}$ then ${\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}={\mathbf{T}}{\mathbf{Z}}{\mathbf{Z}}^{\top}{\mathbf{T}}^{\top}=({\mathbf{T}}{\mathbf{Z}})({\mathbf{T}}{\mathbf{Z}})^{\top}={\mathbf{C}}_{Z}({\mathbf{T}}{\mathbf{Z}})$.
For the second point if we consider $\tilde{L}(a,b)=a\times b$ then we use that the neighbor embedding problem eq. 16 is equivalent to $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}-[{\mathbf{C}}_{X}({\mathbf{X}})]_{ij}\log([{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})=\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}\tilde{L}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij})$ where $[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij}=g({\mathbf{z}}_{i}-{\mathbf{z}}_{j})+\log(R({\mathbf{Z}}))$ with $g=-\log\circ f$. Since $f$ is logarithmically concave $g$ is convex. Moreover we have that $\tilde{L}(a,\cdot)$ is convex (it is linear) and non-decreasing since $a\in\mathbb{R}_{+}$ in this case (${\mathbf{C}}_{\mathbf{X}}$ is non-negative). Also for any ${\mathbf{Z}}\in\mathbb{R}^{N\times d}$ and ${\mathbf{T}}\in\operatorname{DS}$ we have, using Jensen’s inequality since ${\mathbf{T}}$ is doubly-stochastic,
$\begin{split}[{\mathbf{T}}\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij}&=\sum_{kl}g({\mathbf{z}}_{k}-{\mathbf{z}}_{l})T_{ik}T_{jl}+\log(R({\mathbf{Z}}))\\ &\geq g(\sum_{kl}({\mathbf{z}}_{k}-{\mathbf{z}}_{l})T_{ik}T_{jl})+\log(R({\mathbf{Z}}))\\ &=g(\sum_{k}{\mathbf{z}}_{k}T_{ik}-\sum_{l}{\mathbf{z}}_{l}T_{jl})+\log(R({\mathbf{Z}}))\,.\end{split}$ (18)
Now we will prove that $\log(R({\mathbf{Z}}))\geq\log(R({\mathbf{T}}{\mathbf{Z}}))$. Using Birkhoff’s theorem [6] the matrix ${\mathbf{T}}$ can be decomposed as a convex combination of permutation matrices, i.e. , ${\mathbf{T}}=\sum_{k}\lambda_{k}{\mathbf{P}}_{k}$ where $({\mathbf{P}}_{k})_{k}$ are permutation matrices and $\lambda_{k}\in\mathbb{R}_{+}$ with $\sum_{k}\lambda_{k}=1$. Hence by convexity and Jensen’s inequality $R({\mathbf{T}}{\mathbf{Z}})=R(\sum_{k}\lambda_{k}{\mathbf{P}}_{k}{\mathbf{Z}})\leq\sum_{k}\lambda_{k}R({\mathbf{P}}_{k}{\mathbf{Z}})$. Now using that $R$ is permutation invariant we get $R({\mathbf{P}}_{k}{\mathbf{Z}})=R({\mathbf{Z}})$ and thus $R({\mathbf{T}}{\mathbf{Z}})\leq\sum_{k}\lambda_{k}R({\mathbf{Z}})=R({\mathbf{Z}})$. Since the logarithm is non-decreasing we have $\log(R({\mathbf{Z}}))\geq\log(R({\mathbf{T}}{\mathbf{Z}}))$ and, overall,
$\begin{split}[{\mathbf{T}}\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij}\geq g(\sum_{k}{\mathbf{z}}_{k}T_{ik}-\sum_{l}{\mathbf{z}}_{l}T_{jl})+\log(R({\mathbf{T}}{\mathbf{Z}}))=[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{T}}{\mathbf{Z}})]_{ij}\,.\end{split}$ (19)
Thus if we introduce ${\mathbf{Y}}={\mathbf{T}}{\mathbf{Z}}$ we have $[{\mathbf{T}}\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij}\geq[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Y}})]_{ij}$ and eq. 12 is satisfied. Thus we can apply proposition A.2 and state that $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}\tilde{L}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij})$ and $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{\tilde{L}}({\mathbf{C}}_{X}({\mathbf{X}}),\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$ are equivalent which concludes as $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{\tilde{L}}({\mathbf{C}}_{X}({\mathbf{X}}),\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})=\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$. ∎
It remains to prove the second point of theorem 3.2 as stated below.
Proposition A.4.
Consider $\operatorname{Im}({\mathbf{C}}_{X})\subseteq\mathbb{R}_{+}^{N\times N}$, $L=L_{\operatorname{KL}}$. Suppose that for any ${\mathbf{X}}$ the matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD and for any ${\mathbf{Z}}$
${\mathbf{C}}_{Z}({\mathbf{Z}})=\operatorname{diag}(\bm{\alpha}_{\mathbf{Z}}){\mathbf{K}}_{\mathbf{Z}}\operatorname{diag}(\bm{\beta}_{\mathbf{Z}})\,,$ (20)
where $\bm{\alpha}_{\mathbf{Z}},\bm{\beta}_{\mathbf{Z}}\in\mathbb{R}^{N}_{>0}$ and ${\mathbf{K}}_{\mathbf{Z}}\in\mathbb{R}^{N\times N}_{>0}$ is such that $\log({\mathbf{K}}_{\mathbf{Z}})$ is CPD. Then the minimum eq. DR is equal to $\min_{{\mathbf{Z}}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$.
Proof.
To prove this result we will show that, for any ${\mathbf{Z}}$, the function
${\mathbf{T}}\in{\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})\to E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{T}})\,,$ (21)
is actually concave. Indeed, in this case there exists a minimizer which is an extremal point of ${\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$. By Birkhoff’s theorem [6] these extreme points are the matrices $\frac{1}{N}{\mathbf{P}}$ where ${\mathbf{P}}$ is a $N\times N$ permutation matrix. Consequently, when the function eq. 21 is concave minimizing $\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ in ${\mathbf{Z}}$ is equivalent to minimizing in ${\mathbf{Z}}$
$\min_{{\mathbf{P}}\in\mathbb{R}^{N\times N}\text{ permutation}}\ \sum_{ijkl}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ik},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{jl})P_{ij}P_{kl}=\min_{\sigma\in S_{N}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{\sigma(i)\sigma(j)})\,,$ (22)
which is exactly the Gromov-Monge problem described in lemma 3.1 and thus the problem is equivalent to eq. DR by lemma 3.1.
First note that $L(x,y)=x\log(x)-x-x\log(y)+y$ so for any ${\mathbf{T}}\in{\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ the loss $E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{T}})$ is equal to
$\begin{split}\sum_{ijkl}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ik},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{jl})T_{ij}T_{kl}&=a_{\mathbf{X}}+b_{\mathbf{Z}}-\sum_{ijkl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([{\mathbf{C}}_{Z}({\mathbf{Z}})]_{jl})T_{ij}T_{kl}\\ &=a_{\mathbf{X}}+b_{\mathbf{Z}}-\sum_{ijkl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([\bm{\alpha}_{\mathbf{Z}}]_{j}[\bm{\beta}_{\mathbf{Z}}]_{l}[{\mathbf{W}}]_{jl})T_{ij}T_{kl}\\ &=a_{\mathbf{X}}+b_{\mathbf{Z}}-\frac{1}{N}\sum_{ijk}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([\bm{\alpha}_{\mathbf{Z}}]_{j})T_{ij}\\ &-\frac{1}{N}\sum_{ikl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([\bm{\beta}_{\mathbf{Z}}]_{l})T_{kl}\\ &-\sum_{ijkl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([{\mathbf{K}}_{\mathbf{Z}}]_{jl})T_{ij}T_{kl}\,,\end{split}$ (23)
where $a_{\mathbf{X}},b_{\mathbf{Z}}$ are terms that do not depend on ${\mathbf{T}}$. Since the problem is quadratic the concavity only depends on the term $-\sum_{ijkl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([{\mathbf{K}}_{\mathbf{Z}}]_{jl})T_{ij}T_{kl}=-\operatorname{Tr}({\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}^{\top}\log({\mathbf{K}}_{\mathbf{Z}}){\mathbf{T}})$. From [47] we know that the function ${\mathbf{T}}\to-\operatorname{Tr}({\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}^{\top}\log({\mathbf{K}}_{\mathbf{Z}}){\mathbf{T}})$ is concave on ${\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ when ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD and $\log({\mathbf{W}}_{\mathbf{Z}})$ is CPD. This concludes the proof.
A.3 Necessary and sufficient condition
We give a necessary and sufficient condition under which the DR problem is equivalent to a GW problem
Proposition A.5.
Let ${\mathbf{C}}_{1}\in\mathbb{R}^{N\times N},L:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and $\mathcal{C}\subseteq\mathbb{R}^{N\times N}$ a subspace of $N\times N$ matrices. We suppose that $\mathcal{C}$ is stable by permutations i.e. , ${\mathbf{C}}\in\mathcal{C}$ implies that ${\mathbf{P}}{\mathbf{C}}{\mathbf{P}}^{\top}\in\mathcal{C}$ for any $N\times N$ permutation matrix ${\mathbf{P}}$. Then
$\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{(i,j)\in{[\![N]\!]}^{2}}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{ij})=\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\min_{\begin{smallmatrix}{\mathbf{T}}\in\mathbb{R}_{+}^{N\times N}\\ {\mathbf{T}}\bm{1}_{N}=\bm{1}_{N}\\ {\mathbf{T}}^{\top}\bm{1}_{N}=\bm{1}_{N}\end{smallmatrix}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T_{ik}T_{jl}$ (24)
if and only if the optimal assignment problem
$\min_{\sigma_{1},\sigma_{2}\in S_{N}}\ f(\sigma_{1},\sigma_{2}):=\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{\sigma_{1}(i)\sigma_{2}(j)})$ (25)
admits an optimal solution $(\sigma_{1}^{\star},\sigma_{2}^{\star})$ with $\sigma_{1}^{\star}=\sigma_{2}^{\star}$.
Proof.
We note that the LHS of eq. 24 is always smaller than the RHS since ${\mathbf{T}}=\mathbf{I}_{N}$ is admissible for the RHS problem. So both problems are equal if and only if
$\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{(i,j)\in{[\![N]\!]}^{2}}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{ij})\leq\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\min_{\begin{smallmatrix}{\mathbf{T}}\in\mathbb{R}_{+}^{N\times N}\\ {\mathbf{T}}\bm{1}_{N}=\bm{1}_{N}\\ {\mathbf{T}}^{\top}\bm{1}_{N}=\bm{1}_{N}\end{smallmatrix}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T_{ik}T_{jl}\,.$ (26)
Now consider any ${\mathbf{C}}_{2}$ fixed and observe that
$\min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T_{ik}T_{jl}\geq\min_{{\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)}\in\operatorname{DS}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T^{(1)}_{ik}T^{(2)}_{jl}\,.$ (27)
The latter problem is a co-optimal transport problem [60], and, since it is a bilinear problem, there is an optimal solution $({\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)})$ such that both ${\mathbf{T}}^{(1)}$ and ${\mathbf{T}}^{(2)}$ are in an extremal point of $\operatorname{DS}$ which is the space of $N\times N$ permutation matrices by Birkhoff’s theorem [6]. This point was already noted by [37] but we recall the proof for completeness. We note $L_{ijkl}=L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})$ and consider an optimal solution $({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ of $\min_{{\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)}\in\operatorname{DS}}\phi({\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)}):=\sum_{ijkl}L_{ijkl}T^{(1)}_{ik}T^{(2)}_{jl}$. Consider the linear problem $\min_{{\mathbf{T}}\in\operatorname{DS}}\phi({\mathbf{T}},{\mathbf{T}}_{\star}^{(2)})$. Since it is a linear over the space of doubly stochastic matrices it admits a permutation matrix ${\mathbf{P}}^{(1)}$ as optimal solution. Also $\phi({\mathbf{P}}^{(1)},{\mathbf{T}}_{\star}^{(2)})\leq\phi({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ by optimality. Now consider the linear problem $\min_{{\mathbf{T}}\in\operatorname{DS}}\phi({\mathbf{P}}^{(1)},{\mathbf{T}})$, in the same it admits a permutation matrix ${\mathbf{P}}^{(2)}$ as optimal solution and by optimality $\phi({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})\leq\phi({\mathbf{P}}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ thus $\phi({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})\leq\phi({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ which implies that $({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})$ is an optimal solution. Combining with eq. 27 we get
$\begin{split}\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T_{ik}T_{jl}&\geq\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\min_{\sigma_{1},\sigma_{2}\in S_{N}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{\sigma_{1}(i)\sigma_{2}(j)})\\ &=\min_{\sigma_{1},\sigma_{2}\in S_{N}}f(\sigma_{1},\sigma_{2})\,.\end{split}$ (28)
Now suppose that the optimal assignment problem $\min_{\sigma_{1},\sigma_{2}\in S_{N}}\ f(\sigma_{1},\sigma_{2})$ admits an optimal solution $(\sigma_{1}^{\star},\sigma_{2}^{\star})$ with $\sigma_{1}^{\star}=\sigma_{2}^{\star}$. Then with eq. 31
$\begin{split}\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T_{ik}T_{jl}&\geq\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{\sigma^{\star}_{1}(i)\sigma^{\star}_{1}(j)})\\ &\geq\min_{\sigma\in S_{N}}\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{\sigma(i)\sigma(j)})\,.\end{split}$ (29)
Now since $\mathcal{C}$ is stable by permutation then $\{\left([{\mathbf{C}}]_{\sigma(i)\sigma(j)}\right)_{(i,j)\in{[\![N]\!]}^{2}}:{\mathbf{C}}\in\mathcal{C},\sigma\in S_{N}\}=\mathcal{C}$ and consequently $\min_{\sigma\in S_{N}}\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{\sigma(i)\sigma(j)})=\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{ij})$. Consequently, using eq. 29,
$\begin{split}\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T_{ik}T_{jl}&\geq\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{ij})\,,\end{split}$ (30)
and thus both are equal.
Conversely suppose that eq. 24 holds. Then, from eq. 31 we have
$\begin{split}\min_{\sigma_{1},\sigma_{2}\in S_{N}}f(\sigma_{1},\sigma_{2})&=\min_{\sigma_{1},\sigma_{2}\in S_{N}}\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{ij}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{\sigma_{1}(i)\sigma_{2}(j)})\\ &\leq\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ijkl}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})\ T_{ik}T_{jl}\\ &=\min_{{\mathbf{C}}_{2}\in\mathcal{C}}\sum_{(i,j)\in{[\![N]\!]}^{2}}L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{ij})=f(\operatorname{id},\operatorname{id})\,,\end{split}$ (31)
which concludes the proof. ∎
Remark A.6.
The condition on the set of similarity matrices $\mathcal{C}$ is quite reasonable: it indicates that if ${\mathbf{C}}$ is an admissible similarity matrix, then permuting the rows and columns of ${\mathbf{C}}$ results in another admissible similarity matrix. For DR, the corresponding $\mathcal{C}$ is $\mathcal{C}=\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$. In this case, if ${\mathbf{C}}_{Z}({\mathbf{Z}})$ is of the form
$[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}=h(f(\bm{z}_{i},\bm{z}_{j}),g({\mathbf{Z}}))\,,$ (32)
where $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R},\ h:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and $g:\mathbb{R}^{N\times d}\to\mathbb{R}$ which is permutation invariant [9], then $\mathcal{C}$ is stable under permutation. Indeed, permuting the rows and columns of ${\mathbf{C}}_{Z}({\mathbf{Z}})$ by $\sigma$ is equivalent to considering the similarity ${\mathbf{C}}_{Z}({\mathbf{Y}})$, where ${\mathbf{Y}}=(\bm{z}_{\sigma(1)},\cdots,\bm{z}_{\sigma(n)})^{\top}$. Moreover, most similarities in the target space considered in DR take the form eq. 32: $\langle\Phi(\bm{z}_{i}),\Phi(\bm{z}_{j})\rangle_{\mathcal{H}}$ (kernels such as in spectral methods with $\Phi=\operatorname{id}$), $f(\bm{z}_{i},\bm{z}_{j})/\sum_{nm}f(\bm{z}_{n},\bm{z}_{m})$ (normalized similarities such as in SNE and t-SNE). Also note that the condition on $\mathcal{C}=\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$ of proposition A.5 is met as soon as ${\mathbf{C}}_{Z}:\mathbb{R}^{N\times d}\to\mathbb{R}^{N\times N}$ is permutation equivariant.
Appendix B Generalized Semi-relaxed Gromov-Wasserstein is a divergence
Remark B.1 (Weak isomorphism).
According to the notion of weak isomorphism in [17], for a graph $({\mathbf{C}},{\mathbf{h}})$ with corresponding discrete measure $\mu=\sum_{i}h_{i}\delta_{x_{i}}$, two nodes $x_{i}$ and $x_{j}$ are the “same” if they have the same internal perception i.e $C_{ii}=C_{jj}=C_{ij}=C_{ji}$ and external perception $\forall k\neq(i,j),C_{ik}=C_{jk},C_{ki}=C_{kj}$. So two graphs $({\mathbf{C}}_{1},{\mathbf{h}}_{1})$ and $({\mathbf{C}}_{2},{\mathbf{h}}_{2})$ are said to be weakly isomorphic, if there exist a canonical representation $({\mathbf{C}}_{c},{\mathbf{h}}_{c})$ such that $\operatorname{card}(\operatorname{supp}({\mathbf{h}}_{c}))=p\leq n,m$ and $M_{1}\in\left\{0,1\right\}^{n\times p}$ (resp. $M_{2}\in\left\{0,1\right\}^{m\times p}$) such that for $k\in\left\{1,2\right\}$
${\mathbf{C}}_{c}={\mathbf{M}}_{k}^{\top}{\mathbf{C}}_{c}{\mathbf{M}}_{k}\quad\text{and}\quad{\mathbf{h}}_{c}={\mathbf{M}}_{k}^{\top}{\mathbf{h}}_{k}$ (33)
We first emphasize a simple result.
Proposition B.2.
Let any divergence $L:\Omega\times\Omega\rightarrow\mathbb{R}_{+}$ for $\Omega\subseteq\mathbb{R}$, then for any $({\mathbf{C}},{\mathbf{h}})$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$, we have $\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$ if and only if $\operatorname{GW}_{L_{2}}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$.
Proof.
If $\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$, then there exists ${\mathbf{T}}\in\mathcal{U}({\mathbf{h}},\overline{{\mathbf{h}}})$ such that
$E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=\sum_{ijkl}L(C_{ij},\overline{C}_{kl})T_{ik}T_{jl}=0$ (34)
so whenever $T_{ik}T_{jl}\neq 0$, we must have $L(C_{ij},\overline{C}_{kl})=0$ i.e $C_{ij}=\overline{C}_{kl}$ as $L$ is a divergence. Which implies that $E_{L^{\prime}}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=0$ for any other divergence $L^{\prime}$ well defined on any domain $\Omega\times\Omega$, necessarily including $L_{2}$. ∎
Lemma B.3.
Let any divergence $L:\Omega\times\Omega\rightarrow\mathbb{R}_{+}$ for $\Omega\subseteq\mathbb{R}$. Let $({\mathbf{C}},{\mathbf{h}})\in\Omega^{n\times n}\times\Sigma_{n}$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})\in\mathbb{R}^{m\times m}\times\Sigma_{m}$. Then $\operatorname{srGW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$ if and only if there exists $\overline{{\mathbf{h}}}\in\Sigma_{m}$ such that $({\mathbf{C}},{\mathbf{h}})$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ are weakly isomorphic.
Proof.
()($\Rightarrow$)( ⇒ ) As $\operatorname{srGW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$ there exists ${\mathbf{T}}\in\mathcal{U}({\mathbf{h}},\overline{{\mathbf{h}}})$ such that $E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=0$ hence $\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$. Using proposition B.2, $\operatorname{GW}_{L_{2}}=0$ hence using Theorem 18 in [17], it implies that $({\mathbf{C}},{\mathbf{h}})$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ are weakly isomorphic.
()($\Leftarrow$)( ⇐ ) As mentioned, $({\mathbf{C}},{\mathbf{h}})$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ being weakly isomorphic implies that $\operatorname{GW}_{L_{2}}=0$. So there exists ${\mathbf{T}}\in\mathcal{U}({\mathbf{h}},\overline{{\mathbf{h}}})$, such that $E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=0$. Moreover ${\mathbf{T}}$ is admissible for the srGW problem as ${\mathbf{T}}\in\mathcal{U}({\mathbf{h}},\overline{{\mathbf{h}}})\subset\mathcal{U}_{n}({\mathbf{h}})$, thus $\operatorname{srGW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$. ∎
B.1 About trivial solutions of semi-relaxed GW when $L$ is not a proper divergence
We briefly describe here some trivial solutions of $\operatorname{srGW}_{L}$ when $L$ is not a proper divergence. We recall that
$\operatorname{srGW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}})=\min_{{\mathbf{T}}\in\mathbb{R}_{+}^{N\times n}:{\mathbf{T}}\bm{1}_{n}={\mathbf{h}}}E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=\sum_{ijkl}L(C_{ij},\overline{C}_{kl})T_{ik}T_{jl}\,.$ (35)
Suppose that ${\mathbf{h}}\in\Sigma_{N}^{*}$ and that $\overline{{\mathbf{C}}}$ has a minimum value on its diagonal i.e.  $\min_{(i,j)}\overline{C}_{ij}=\min_{ii}\overline{C}_{ii}$. Suppose also that $\forall a,L(a,\cdot)$ is both convex and non-decreasing. First we have $\sum_{j}\frac{T_{ij}}{h_{i}}=1$ for any $i\in{[\![N]\!]}$. Hence using the convexity of $L$, Jensen inequality and the fact that $L(a,\cdot)$ is non-decreasing for any $a$
$\begin{split}\sum_{ijkl}L(C_{ij},\overline{C}_{kl})T_{ik}T_{jl}&=\sum_{ijkl}L(C_{ij},\overline{C}_{kl})h_{i}h_{j}\frac{T_{ik}}{h_{i}}\frac{T_{jl}}{h_{j}}\\ &\geq\sum_{ij}L(C_{ij},\sum_{kl}\overline{C}_{kl}\frac{T_{ik}}{h_{i}}\frac{T_{jl}}{h_{j}})h_{i}h_{j}\\ &\geq\sum_{ij}L(C_{ij},\sum_{kl}(\min_{nm}\overline{C}_{nm})\frac{T_{ik}}{h_{i}}\frac{T_{jl}}{h_{j}})h_{i}h_{j}\\ &=\sum_{ij}L(C_{ij},(\min_{nm}\overline{C}_{nm}))h_{i}h_{j}\\ &=\sum_{ij}L(C_{ij},(\min_{nn}\overline{C}_{nn}))h_{i}h_{j}\,.\end{split}$ (36)
Now suppose without loss of generality that $\min_{ii}\overline{C}_{ii}=\overline{C}_{11}$ then this gives
$\min_{{\mathbf{T}}\in\mathbb{R}_{+}^{N\times n}:{\mathbf{T}}\bm{1}_{n}={\mathbf{h}}}\sum_{ijkl}L(C_{ij},\overline{C}_{kl})T_{ik}T_{jl}\geq\sum_{ij}L(C_{ij},\overline{C}_{11})h_{i}h_{j}\,.$ (37)
Now consider the coupling ${\mathbf{T}}^{\star}=\begin{pmatrix}h_{1}&0&0&0\\ h_{2}&0&0&0\\ \vdots&\vdots&\vdots&\vdots\\ h_{N}&0&0&0\\ \end{pmatrix}$. It is admissible and satisfies
$E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}}^{\star})=\sum_{ij}L(C_{ij},\overline{C}_{11})h_{i}h_{j}\leq\min_{{\mathbf{T}}\in\mathbb{R}_{+}^{N\times n}:{\mathbf{T}}\bm{1}_{n}={\mathbf{h}}}E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})\,.$ (38)
Consequently the coupling ${\mathbf{T}}^{\star}$ is optimal. However the solution given by this coupling is trivial: it consists in sending all the mass to one unique point. In another words, all the nodes in the input graph are sent to a unique node in the target graph. Note that this phenomena is impossible for standard GW because of the coupling constraints.
We emphasize that this hypothesis on $L$ cannot be satisfied as soon as $L$ is a proper divergence. Indeed when $L$ is a divergence the constraint “$L(a,\cdot)$ is non-decreasing for any $a$” is not possible as it would break the divergence constraints $\forall a,b\ L(a,b)\geq 0\text{ and }L(a,b)=0\iff a=b$ (at some point $L$ must be decreasing).
Appendix C Clustering properties: Proof of theorem 4.1
We recall that a matrix ${\mathbf{C}}\in\mathbb{R}^{N\times N}$ is conditionally positive definite (CPD), resp. negative definite (CND), if $\forall{\mathbf{x}}\in\mathbb{R}^{N},{\mathbf{x}}^{\top}\mathbf{1}_{N}=0\text{s.t. }{\mathbf{x}}^{\top}{\mathbf{C}}{\mathbf{x}}\geq 0$, resp. $\leq 0$. We also consider the Hadamard product of matrices as ${\mathbf{A}}\odot{\mathbf{B}}=(A_{ij}\times B_{ij})_{ij}$. The $i$-th column of a matrix ${\mathbf{T}}$ is the vector denoted by ${\mathbf{T}}_{:,i}$. For a vector ${\mathbf{x}}\in\mathbb{R}^{n}$ we denote by $\operatorname{diag}({\mathbf{x}})$ the diagonal $n\times n$ matrix whose elements are the $x_{i}$.
We state below the theorem that we prove in this section. See 4.1 In order to prove this result we introduce the space of semi-relaxed couplings whose columns are not zero
${\mathcal{U}}_{n}^{+}({\mathbf{h}}_{X})=\{{\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X}):\forall i\in{[\![n]\!]},{\mathbf{T}}_{:,i}\neq 0\}\,,$ (39)
and we will use the following lemma.
Lemma C.1.
Let ${\mathbf{h}}_{X}\in\Sigma_{N}$, $L=L_{2}$ and ${\mathbf{C}}_{X}({\mathbf{X}})$ symmetric. For any ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})$, the matrix $\overline{{\mathbf{C}}}({\mathbf{T}})\in\mathbb{R}^{n\times n}$ defined by
$\overline{{\mathbf{C}}}({\mathbf{T}})=\begin{cases}{\mathbf{T}}_{:,i}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}_{:,j}/({\mathbf{T}}_{:,i}^{\top}\bm{1}_{N})({\mathbf{T}}_{:,j}^{\top}\bm{1}_{N})&\text{ for }(i,j)\text{ such that}{\mathbf{T}}_{:,i}\text{ and }{\mathbf{T}}_{:,j}\neq 0\\ 0&\text{ otherwise}\end{cases}$ (40)
is a minimizer of $G:\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}\to E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$. For ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})$ the expression of the minimum is
$G({\mathbf{T}})=\operatorname{cte}-\operatorname{Tr}\left(({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}(\overline{{\mathbf{C}}}({\mathbf{T}})\odot\overline{{\mathbf{C}}}({\mathbf{T}}))\right)\,,$ (41)
which defines a continuous function on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$. If ${\mathbf{T}}\in{\mathcal{U}}_{n}^{+}({\mathbf{h}}_{X})$ it becomes
$G({\mathbf{T}})=\operatorname{cte}-\sum_{ij}\frac{({\mathbf{T}}_{:,i}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}_{:,j})^{2}}{({\mathbf{T}}_{:,i}^{\top}\bm{1}_{N})({\mathbf{T}}_{:,j}^{\top}\bm{1}_{N})}=\operatorname{cte}-\|\operatorname{diag}({\mathbf{T}}^{\top}\bm{1}_{N})^{-\frac{1}{2}}{\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\operatorname{diag}({\mathbf{T}}^{\top}\bm{1}_{N})^{-\frac{1}{2}}\|_{F}^{2}\,.$ (42)
Proof.
First see that $\overline{{\mathbf{C}}}({\mathbf{T}})$ is well defined since ${\mathbf{T}}_{:,i}\neq 0\iff{\mathbf{T}}_{:,i}^{\top}\bm{1}_{N}\neq 0$ because ${\mathbf{T}}$ is non-negative. Consider, for ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})$, the function
$F({\mathbf{T}},\overline{{\mathbf{C}}}):=E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})=\sum_{ijkl}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}-\overline{C}_{jl})^{2}T_{ij}T_{kl}\,,$ (43)
A development yields (using ${\mathbf{T}}^{\top}\bm{1}_{N}={\mathbf{h}}_{X}$)
$\begin{split}F({\mathbf{T}},\overline{{\mathbf{C}}})&=\sum_{ik}[{\mathbf{C}}_{X}({\mathbf{X}})]^{2}_{ik}[{\mathbf{h}}_{X}]_{i}[{\mathbf{h}}_{X}]_{k}+\sum_{jl}\overline{C}^{2}_{jl}(\sum_{i}T_{ij})(\sum_{k}T_{kl})-2\sum_{ijkl}\overline{C}_{jl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}T_{ij}T_{kl}\,.\end{split}$ (44)
We can rewrite $\sum_{ijkl}\overline{C}_{jl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}T_{ij}T_{kl}=\operatorname{Tr}({\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{{\mathbf{C}}})$. Also we have $(\sum_{i}T_{ij})(\sum_{k}T_{kl})=[({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}]_{jl}$ Thus
$\begin{split}\sum_{jl}\overline{C}^{2}_{jl}(\sum_{i}T_{ij})(\sum_{k}T_{kl})&=\operatorname{Tr}(({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}(\overline{{\mathbf{C}}}\odot\overline{{\mathbf{C}}}))\,.\\ \end{split}$ (45)
Overall
$F({\mathbf{T}},\overline{{\mathbf{C}}})=\operatorname{cte}+\operatorname{Tr}(({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}(\overline{{\mathbf{C}}}\odot\overline{{\mathbf{C}}}))-2\operatorname{Tr}({\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{\mathbf{C}})\,.$ (46)
Now taking the derivative with respect to $\overline{\mathbf{C}}$, the first order conditions are
$\partial_{2}F({\mathbf{T}},\overline{{\mathbf{C}}})=2(\overline{\mathbf{C}}\odot({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}-{\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}})=0\,.$ (47)
For $(i,j)$ such that ${\mathbf{T}}_{:,i}\text{ and }{\mathbf{T}}_{:,j}\neq 0$ we have $[\partial_{2}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))]_{ij}=0$. For $(i,j)$ such that ${\mathbf{T}}_{:,i}\text{ or }{\mathbf{T}}_{:,j}=0$ we have $[\overline{\mathbf{C}}\odot({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}]_{ij}=0$ and also $[{\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}]_{ij}={\mathbf{T}}_{:,i}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}_{:,j}=0$. In particular the matrix $\overline{{\mathbf{C}}}({\mathbf{T}})$ satisfies the first order conditions. When $L=L_{2}$ the problem $\min_{\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})=\frac{1}{2}\sum_{ijkl}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}-\overline{C}_{jl})^{2}T_{ij}T_{kl}$ is convex in $\overline{{\mathbf{C}}}$. The first order conditions are sufficient hence $\overline{{\mathbf{C}}}({\mathbf{T}})$ is a minimizer.
Also ${\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}=\overline{{\mathbf{C}}}({\mathbf{T}})\odot({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}$ by definition of $\overline{{\mathbf{C}}}({\mathbf{T}})$ thus
$\begin{split}\operatorname{Tr}({\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{\mathbf{C}}({\mathbf{T}}))&=\operatorname{Tr}([\overline{{\mathbf{C}}}({\mathbf{T}})\odot({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}]\overline{\mathbf{C}}({\mathbf{T}}))\\ &=\operatorname{Tr}(({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}[\overline{{\mathbf{C}}}({\mathbf{T}})\odot\overline{{\mathbf{C}}}({\mathbf{T}})])\,.\end{split}$ (48)
Hence
$\begin{split}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))&=\operatorname{cte}-\operatorname{Tr}\left(({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}(\overline{{\mathbf{C}}}({\mathbf{T}})\odot\overline{{\mathbf{C}}}({\mathbf{T}}))\right)\,.\end{split}$ (49)
Consequently for ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})$ such that $\forall i\in{[\![n]\!]},{\mathbf{T}}_{:,i}\neq 0$ we have
$\begin{split}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))&=\operatorname{cte}-\sum_{ij}\frac{({\mathbf{T}}_{:,i}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}_{:,j})^{2}}{({\mathbf{T}}_{:,i}^{\top}\bm{1}_{N})({\mathbf{T}}_{:,j}^{\top}\bm{1}_{N})}\\ &=\operatorname{cte}-\|\operatorname{diag}({\mathbf{T}}^{\top}\bm{1}_{N})^{-\frac{1}{2}}{\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\operatorname{diag}({\mathbf{T}}^{\top}\bm{1}_{N})^{-\frac{1}{2}}\|_{F}^{2}\,.\end{split}$ (50)
It just remains to demonstrate the continuity of $G$. We consider for $({\mathbf{x}},{\mathbf{y}})\in\mathbb{R}_{+}^{N}\times\mathbb{R}_{+}^{N}$ the function
$g({\mathbf{x}},{\mathbf{y}})=\begin{cases}\frac{({\mathbf{x}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{y}})^{2}}{\|{\mathbf{x}}\|_{1}\|{\mathbf{y}}\|_{1}}&\text{ when }{\mathbf{x}}\neq 0\text{ and }{\mathbf{y}}\neq 0\\ 0&\text{ otherwise}\end{cases}$ (51)
and we show that $g$ is continuous. For $({\mathbf{x}},{\mathbf{y}})\neq(0,0)$ this is clear. Now using that
$0\leq g({\mathbf{x}},{\mathbf{y}})=\frac{(\sum_{ij}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ij}x_{i}y_{j})^{2}}{(\sum_{i}x_{i})(\sum_{j}y_{j})}\leq\|{\mathbf{C}}_{X}({\mathbf{X}})\|^{2}_{\infty}\frac{(\sum_{i}x_{i})^{2}(\sum_{j}y_{j})^{2}}{(\sum_{i}x_{i})(\sum_{j}y_{j})}=\|{\mathbf{C}}_{X}({\mathbf{X}})\|^{2}_{\infty}\|{\mathbf{x}}\|_{1}\|{\mathbf{y}}\|_{1}\,,$ (52)
this shows $\lim_{{\mathbf{x}}\to 0}g({\mathbf{x}},{\mathbf{y}})=0=g(0,{\mathbf{y}})$ and $\lim_{{\mathbf{y}}\to 0}g({\mathbf{x}},{\mathbf{y}})=0=g({\mathbf{x}},0)$. Now for ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})$ we have $G({\mathbf{T}})=\operatorname{cte}-\sum_{ij}g({\mathbf{T}}_{:,i},{\mathbf{T}}_{:,j})$ which defines a continuous function.
To prove the theorem we will first prove that the function $G:{\mathbf{T}}\to\min_{\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ is concave on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$ and by a continuity argument it will be concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$. The concavity will allow us to prove that the minimum of $G$ is achieved in an extreme point of ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$ which is a membership matrix.
Proposition C.2.
Let ${\mathbf{h}}_{\mathbf{X}}\in\Sigma_{N},L=L_{2}$ and suppose that ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD or CND. Then the function $G:{\mathbf{T}}\to\min_{\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ is concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$. Consequently theorem 4.1 holds.
Proof.
We recall that $F({\mathbf{T}},\overline{{\mathbf{C}}}):=E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ and $G({\mathbf{T}})=F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))$. From lemma C.1 we know that $\overline{{\mathbf{C}}}({\mathbf{T}})$ is a minimizer of $\overline{{\mathbf{C}}}\to F({\mathbf{T}},\overline{{\mathbf{C}}})$ hence it satisfies the first order conditions $\partial_{2}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))=0$. Every quantity is differentiable on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$. Hence, taking the derivative of $G$ and using the first order conditions
$\nabla G({\mathbf{T}})=\partial_{1}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))+\partial_{2}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))[\nabla\overline{{\mathbf{C}}}({\mathbf{T}})]=\partial_{1}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))\,.$ (53)
We will found the expression of this gradient. In the proof of lemma C.1 we have seen that
$\begin{split}F({\mathbf{T}},\overline{{\mathbf{C}}})&=\operatorname{cte}+\operatorname{Tr}(({\mathbf{T}}^{\top}\bm{1}_{N})({\mathbf{T}}^{\top}\bm{1}_{N})^{\top}(\overline{{\mathbf{C}}}\odot\overline{{\mathbf{C}}}))-2\operatorname{Tr}({\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{\mathbf{C}})\\ &=\operatorname{cte}+\operatorname{Tr}({\mathbf{T}}^{\top}\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{T}}(\overline{{\mathbf{C}}}\odot\overline{{\mathbf{C}}}))-2\operatorname{Tr}({\mathbf{T}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{\mathbf{C}})\,.\end{split}$ (54)
Using that the derivative of ${\mathbf{T}}\to\operatorname{Tr}({\mathbf{T}}^{\top}{\mathbf{A}}{\mathbf{T}}{\mathbf{B}})$ is ${\mathbf{A}}^{\top}{\mathbf{T}}{\mathbf{B}}^{\top}+{\mathbf{A}}{\mathbf{T}}{\mathbf{B}}$ and that ${\mathbf{C}}_{X}({\mathbf{X}})$ is symmetric we get
$\partial_{1}F({\mathbf{T}},\overline{{\mathbf{C}}})=\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{T}}(\overline{{\mathbf{C}}}\odot\overline{{\mathbf{C}}})^{\top}+\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{T}}(\overline{{\mathbf{C}}}\odot\overline{{\mathbf{C}}})-2{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{{\mathbf{C}}}^{\top}-2{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{{\mathbf{C}}}\,.$ (55)
Finally, applying to the symmetric matrix $\overline{{\mathbf{C}}}=\overline{{\mathbf{C}}}({\mathbf{T}})$
$\nabla G({\mathbf{T}})=\partial_{1}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))=2\left(\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{T}}(\overline{{\mathbf{C}}}({\mathbf{T}})\odot\overline{{\mathbf{C}}}({\mathbf{T}}))-2{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}\overline{{\mathbf{C}}}({\mathbf{T}})\right)\,.$ (56)
In what follows we define
$D({\mathbf{T}}):=\operatorname{diag}({\mathbf{T}}^{\top}\bm{1}_{N})^{-1}\in\mathbb{R}^{n\times n}\,,$ (57)
when applicable. Using the expression of the gradient we will show that $G$ is concave on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$ and we will conclude by a continuity argument on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$. Take $({\mathbf{P}},{\mathbf{Q}})\in{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})\times{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$ we will prove
$G({\mathbf{P}})-G({\mathbf{Q}})-\langle\nabla G({\mathbf{Q}}),{\mathbf{P}}-{\mathbf{Q}}\rangle\leq 0\,.$ (58)
From lemma C.1 we have the expression (since ${\mathbf{P}}\in{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$)
$G({\mathbf{P}})=\operatorname{cte}-\|D({\mathbf{P}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{P}}D({\mathbf{P}})^{\frac{1}{2}}\|_{F}^{2}\,,$ (59)
(same for $G({\mathbf{Q}})$) and
$\overline{{\mathbf{C}}}({\mathbf{Q}})=D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})\,.$ (60)
We will now calculate $\langle\nabla G({\mathbf{Q}}),{\mathbf{P}}\rangle$ which involves $\langle\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{Q}}(\overline{{\mathbf{C}}}({\mathbf{Q}})\odot\overline{{\mathbf{C}}}({\mathbf{Q}}),{\mathbf{P}}\rangle$ and $\langle{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}\overline{{\mathbf{C}}}({\mathbf{Q}}),{\mathbf{P}}\rangle$. For the first term we have
$\begin{split}&\langle\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{Q}}(\overline{{\mathbf{C}}}({\mathbf{Q}})\odot\overline{{\mathbf{C}}}({\mathbf{Q}})),{\mathbf{P}}\rangle\\ &=\operatorname{Tr}({\mathbf{P}}^{\top}\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{Q}}\overline{{\mathbf{C}}}({\mathbf{Q}})^{\odot 2})=\operatorname{Tr}(\bm{1}_{N}^{\top}{\mathbf{Q}}\overline{{\mathbf{C}}}({\mathbf{Q}})^{\odot 2}{\mathbf{P}}^{\top}\bm{1}_{N})\\ &=({\mathbf{Q}}^{\top}\bm{1}_{N})^{\top}(\overline{{\mathbf{C}}}({\mathbf{Q}})\odot\overline{{\mathbf{C}}}({\mathbf{Q}})){\mathbf{P}}^{\top}\bm{1}_{N}\\ &=\operatorname{Tr}(\overline{{\mathbf{C}}}({\mathbf{Q}})\operatorname{diag}({\mathbf{Q}}^{\top}\bm{1}_{N})\overline{{\mathbf{C}}}({\mathbf{Q}})\operatorname{diag}({\mathbf{P}}^{\top}\bm{1}_{N}))\\ &=\operatorname{Tr}\left([D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})]D({\mathbf{Q}})^{-1}[D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})]D({\mathbf{P}})^{-1}\right)\\ &=\operatorname{Tr}(D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})D({\mathbf{P}})^{-1})\\ &=\operatorname{Tr}(D({\mathbf{P}})^{-\frac{1}{2}}[D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})]D({\mathbf{P}})^{-\frac{1}{2}})\\ &=\operatorname{Tr}(D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}D({\mathbf{Q}})^{\frac{1}{2}}{\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})D({\mathbf{P}})^{-\frac{1}{2}})\\ &=\langle D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}},D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\rangle\\ &=\|D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\|_{F}^{2}\,.\end{split}$ (61)
For the second term
$\begin{split}&\langle{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}\overline{{\mathbf{C}}}({\mathbf{Q}}),{\mathbf{P}}\rangle\\ &=\operatorname{Tr}({\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}}))\\ &=\operatorname{Tr}(D({\mathbf{Q}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}D({\mathbf{Q}})^{\frac{1}{2}}{\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}})\\ &=\langle D({\mathbf{Q}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}},D({\mathbf{Q}})^{\frac{1}{2}}{\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\rangle\\ &=\langle D({\mathbf{Q}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}},D({\mathbf{P}})^{\frac{1}{2}}D({\mathbf{Q}})^{-\frac{1}{2}}D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\rangle\\ &=\langle D({\mathbf{P}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}},D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\rangle\,.\end{split}$ (62)
This gives
$\begin{split}\langle\nabla G({\mathbf{Q}}),{\mathbf{P}}\rangle&=2\|D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\|_{F}^{2}\\ &-4\langle D({\mathbf{P}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}},D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\rangle\\ &=2\|D({\mathbf{P}})^{-\frac{1}{2}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}-D({\mathbf{P}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\|_{F}^{2}\\ &-2\|D({\mathbf{P}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\|_{F}^{2}\,.\\ \end{split}$ (63)
From this equation we get directly that
$\begin{split}\langle\nabla G({\mathbf{Q}}),{\mathbf{Q}}\rangle&=-2\|D({\mathbf{Q}})^{\frac{1}{2}}{\mathbf{Q}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\|_{F}^{2}\\ \text{ and }\langle\nabla G({\mathbf{Q}}),{\mathbf{P}}\rangle&\geq-2\|D({\mathbf{P}})^{\frac{1}{2}}{\mathbf{P}}^{\top}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{Q}}D({\mathbf{Q}})^{\frac{1}{2}}\|_{F}^{2}\,.\ \end{split}$ (64)
Hence
[…]
We note ${\mathbf{U}}={\mathbf{P}}D({\mathbf{P}}){\mathbf{P}}^{\top}\in\mathbb{R}^{N\times N},{\mathbf{V}}={\mathbf{Q}}D({\mathbf{Q}}){\mathbf{Q}}^{\top}\in\mathbb{R}^{N\times N}$, the previous calculus shows that
$\begin{split}G({\mathbf{P}})-G({\mathbf{Q}})-\langle\nabla G({\mathbf{Q}}),{\mathbf{P}}-{\mathbf{Q}}\rangle&\leq-\operatorname{Tr}({\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))-\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}))\\ &+2\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))\,,\end{split}$ (66)
Now note that
${\mathbf{U}}^{\top}\bm{1}_{N}={\mathbf{P}}D({\mathbf{P}}){\mathbf{P}}^{\top}\bm{1}_{N}={\mathbf{P}}\operatorname{diag}({\mathbf{P}}^{\top}\bm{1}_{N})^{-1}{\mathbf{P}}^{\top}\bm{1}_{N}={\mathbf{P}}\bm{1}_{N}={\mathbf{h}}_{X}\,.$ (67)
Since ${\mathbf{U}}$ is symmetric we also have ${\mathbf{U}}\bm{1}_{N}={\mathbf{h}}_{X}$ and similarly we have the same result for ${\mathbf{V}}$. Overall ${\mathbf{V}}^{\top}\bm{1}_{N}={\mathbf{U}}^{\top}\bm{1}_{N}$ and ${\mathbf{V}}\bm{1}_{N}={\mathbf{U}}\bm{1}_{N}$. Since ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD or CND we can apply lemma C.3 below which proves that $-\operatorname{Tr}({\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))-\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}))+2\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))\leq 0$ and consequently that $G$ is concave on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$. We now use the continuity of $G$ to prove that it is concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$.
Take ${\mathbf{P}}\in{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$ and ${\mathbf{Q}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})\setminus{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$ i.e.  there exists $k\in{[\![n]\!]}$ such that ${\mathbf{Q}}_{:,k}=0$. Without loss of generality we suppose $k=1$. Consider for $m\in\mathbb{N}^{*}$ the matrix ${\mathbf{Q}}^{(m)}=(\frac{1}{m}\bm{1}_{N},{\mathbf{Q}}_{:,2},\cdots,{\mathbf{Q}}_{:,n})$. Then ${\mathbf{Q}}^{(m)}\to{\mathbf{Q}}$ as $m\to+\infty$. Also since ${\mathbf{Q}}^{(m)}\in{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$ we have by concavity of $G$
$G((1-\lambda){\mathbf{P}}+\lambda{\mathbf{Q}}^{(m)})\geq(1-\lambda)G({\mathbf{P}})+\lambda G({\mathbf{Q}}^{(m)})\,,$ (68)
for any $\lambda\in[0,1]$. Taking the limit as $m\to\infty$ gives, by continuity of $G$,
$G((1-\lambda){\mathbf{P}}+\lambda{\mathbf{Q}})\geq(1-\lambda)G({\mathbf{P}})+\lambda G({\mathbf{Q}})\,,$ (69)
and hence $G$ is concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$. This proves theorem 4.1. Indeed the minimization of ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})\to\min_{\overline{{\mathbf{C}}}}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ is a minimization of a concave function over a polytope, hence admits an extremity of ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$ as minimizer. But these extremities are membership matrices as they can be described as $\{\operatorname{diag}({\mathbf{h}}_{X}){\mathbf{P}}:{\mathbf{P}}\in\{0,1\}^{N\times n},{\mathbf{P}}^{\top}\bm{1}_{n}=\bm{1}_{N}\}$ [11]. ∎
Lemma C.3.
Let ${\mathbf{C}}\in\mathbb{R}^{N\times N}$ be a CPD or CND matrix. Then for any $({\mathbf{P}},{\mathbf{Q}})\in\mathbb{R}^{N\times N}\times\mathbb{R}^{N\times N}$ such that ${\mathbf{P}}^{\top}\bm{1}_{N}={\mathbf{Q}}^{\top}\bm{1}_{N}$ and ${\mathbf{P}}\bm{1}_{N}={\mathbf{Q}}\bm{1}_{N}$ we have
$\operatorname{Tr}({\mathbf{P}}^{\top}{\mathbf{C}}{\mathbf{Q}}{\mathbf{C}})\leq\frac{1}{2}(\operatorname{Tr}({\mathbf{P}}^{\top}{\mathbf{C}}{\mathbf{P}}{\mathbf{C}})+\operatorname{Tr}({\mathbf{Q}}^{\top}{\mathbf{C}}{\mathbf{Q}}{\mathbf{C}}))\,.$ (70)
Proof.
First, since ${\mathbf{C}}$ is symmetric,
$\begin{split}\operatorname{Tr}\left(({\mathbf{P}}-{\mathbf{Q}})^{\top}{\mathbf{C}}({\mathbf{P}}-{\mathbf{Q}}){\mathbf{C}}\right)&=\operatorname{Tr}({\mathbf{P}}^{\top}{\mathbf{C}}{\mathbf{P}}{\mathbf{C}}-{\mathbf{P}}^{\top}{\mathbf{C}}{\mathbf{Q}}{\mathbf{C}}-{\mathbf{Q}}^{\top}{\mathbf{C}}{\mathbf{P}}{\mathbf{C}}+{\mathbf{Q}}^{\top}{\mathbf{C}}{\mathbf{Q}}{\mathbf{C}})\\ &=\operatorname{Tr}({\mathbf{P}}^{\top}{\mathbf{C}}{\mathbf{P}}{\mathbf{C}})+\operatorname{Tr}({\mathbf{Q}}^{\top}{\mathbf{C}}{\mathbf{Q}}{\mathbf{C}})-2\operatorname{Tr}({\mathbf{P}}^{\top}{\mathbf{C}}{\mathbf{Q}}{\mathbf{C}})\,.\end{split}$ (71)
We note ${\mathbf{U}}={\mathbf{P}}-{\mathbf{Q}}$. Since ${\mathbf{P}}^{\top}\bm{1}_{N}={\mathbf{Q}}^{\top}\bm{1}_{N}$ we have ${\mathbf{U}}^{\top}\bm{1}_{N}=0$. In the same way ${\mathbf{U}}\bm{1}_{N}=0$. We introduce ${\mathbf{H}}={\mathbf{I}}_{N}-\frac{1}{N}\bm{1}_{N}\bm{1}_{N}^{\top}$ the centering matrix. Note that
${\mathbf{H}}{\mathbf{U}}{\mathbf{H}}=({\mathbf{U}}-\frac{1}{N}\bm{1}_{N}(\bm{1}_{N}^{\top}{\mathbf{U}})){\mathbf{H}}={\mathbf{U}}{\mathbf{H}}={\mathbf{U}}-\frac{1}{N}({\mathbf{U}}\bm{1}_{N})\bm{1}_{N}^{\top}={\mathbf{U}}\,.$ (72)
Also ${\mathbf{C}}$ is CPD if and only if ${\mathbf{H}}{\mathbf{C}}{\mathbf{H}}$ is positive semi-definite (PSD). Indeed if ${\mathbf{H}}{\mathbf{C}}{\mathbf{H}}$ is PSD then take ${\mathbf{x}}$ such that ${\mathbf{x}}^{\top}\bm{1}_{N}=0$. We then have ${\mathbf{H}}{\mathbf{x}}={\mathbf{x}}$ and thus ${\mathbf{x}}^{\top}{\mathbf{C}}{\mathbf{x}}={\mathbf{x}}^{\top}({\mathbf{H}}{\mathbf{C}}{\mathbf{H}}){\mathbf{x}}\geq 0$. On the other hand when ${\mathbf{C}}$ is CPD then take any ${\mathbf{x}}$ and see that ${\mathbf{x}}^{\top}{\mathbf{H}}{\mathbf{C}}{\mathbf{H}}{\mathbf{x}}=({\mathbf{H}}{\mathbf{x}})^{\top}{\mathbf{C}}({\mathbf{H}}{\mathbf{x}})$. But $({\mathbf{H}}{\mathbf{x}})^{\top}\bm{1}_{N}={\mathbf{x}}^{\top}({\mathbf{H}}^{\top}\bm{1}_{N})=0$. So $({\mathbf{H}}{\mathbf{x}})^{\top}{\mathbf{C}}({\mathbf{H}}{\mathbf{x}})\geq 0$.
By hypothesis ${\mathbf{C}}$ is CPD so ${\mathbf{H}}{\mathbf{C}}{\mathbf{H}}$ is PSD and symmetric, so it has a square root. But using eq. 72 we get
$\begin{split}\operatorname{Tr}\left(({\mathbf{P}}-{\mathbf{Q}})^{\top}{\mathbf{C}}({\mathbf{P}}-{\mathbf{Q}}){\mathbf{C}}\right)&=\operatorname{Tr}({\mathbf{U}}^{\top}{\mathbf{C}}{\mathbf{U}}{\mathbf{C}})=\operatorname{Tr}({\mathbf{H}}{\mathbf{U}}^{\top}{\mathbf{H}}{\mathbf{C}}{\mathbf{H}}{\mathbf{U}}{\mathbf{H}}{\mathbf{C}})\\ &=\operatorname{Tr}({\mathbf{U}}^{\top}({\mathbf{H}}{\mathbf{C}}{\mathbf{H}}){\mathbf{U}}({\mathbf{H}}{\mathbf{C}}{\mathbf{H}}))\\ &=\|({\mathbf{H}}{\mathbf{C}}{\mathbf{H}})^{\frac{1}{2}}{\mathbf{U}}({\mathbf{H}}{\mathbf{C}}{\mathbf{H}})^{\frac{1}{2}}\|_{F}^{2}\geq 0\,,\end{split}$ (73)
For the CND case is suffices to use that ${\mathbf{C}}$ is CND if and only if $-{\mathbf{C}}$ is CPD and that $\operatorname{Tr}\left(({\mathbf{P}}-{\mathbf{Q}})^{\top}{\mathbf{C}}({\mathbf{P}}-{\mathbf{Q}}){\mathbf{C}}\right)=\operatorname{Tr}\left(({\mathbf{P}}-{\mathbf{Q}})^{\top}(-{\mathbf{C}})({\mathbf{P}}-{\mathbf{Q}})(-{\mathbf{C}})\right)$ which concludes the proof. ∎
Appendix D Algorithmic details
We detail in the following the algorithms mentioned in Section 4 to address the semi-relaxed GW divergence computation in our Block Coordinate Descent algorithm for the DistR problem. We begin with details on the computation of an equivalent objective function and its gradient, potentially under low-rank assumptions over structures ${\mathbf{C}}$ and $\overline{{\mathbf{C}}}$.
D.1 Objective function and gradient computation.
Problem statement.
Let consider any matrices ${\mathbf{C}}\in\mathbb{R}^{n\times n}$, $\overline{{\mathbf{C}}}\in\mathbb{R}^{m\times m}$, and a probability vector ${\mathbf{h}}\in\Sigma_{n}$. In all our use cases, we considered inner losses $L:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ which can be decomposed following Proposition 1 in [58]. Namely we assume the existence of functions $f_{1},f_{2},h_{1},h_{2}$ such that
$\forall a,b\in\Omega^{2},\quad L(a,b)=f_{1}(a)+f_{2}(b)-h_{1}(a)h_{2}(b)$ (74)
More specifically we considered
$\begin{split}L_{2}(a,b)=(a-b)^{2}&\implies f_{1}(a)=a^{2},\quad f_{2}(b)=b^{2},\quad h_{1}(a)=a,\quad h_{2}(b)=2b\\ L_{KL}(a,b)=a\log\frac{a}{b}-a+b&\implies f_{1}(a)=a\log a-a,\quad f_{2}(b)=b,\quad h_{1}(a)=a,\quad h_{2}(b)=\log b\\ \end{split}$ (L2)
In this setting, we proposed to solve for the equivalent problem to $\operatorname{srGW}_{L}$ :
$\min_{{\mathbf{T}}\in\mathcal{U}_{n}({\mathbf{h}})}F({\mathbf{T}})$ (srGW-2)
where the objective function reads as
$\begin{split}F({\mathbf{T}})&:=\langle F_{1}(\overline{{\mathbf{C}}},{\mathbf{T}})-F_{2}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}}),{\mathbf{T}}\rangle\\ &=\langle\bm{1}_{N}\bm{1}_{N}^{\top}{\mathbf{T}}f_{2}(\overline{{\mathbf{C}}}),{\mathbf{T}}\rangle-\langle h_{1}({\mathbf{C}}){\mathbf{T}}h_{2}(\overline{{\mathbf{C}}})^{\top},{\mathbf{T}}\rangle\end{split}$ (75)
Problem srGW-2 is usually a non-convex QP with Hessian $\mathcal{H}=f_{2}(\overline{{\mathbf{C}}})\otimes\bm{1}\bm{1}^{\top}-h_{2}(\overline{{\mathbf{C}}})\otimes_{K}h_{1}({\mathbf{C}})$. In all cases this equivalent form is interesting as it avoids computing the constant term $\langle f_{1}({\mathbf{C}}),{\mathbf{h}}{\mathbf{h}}^{\top}\rangle$ that requires $O(N^{2})$ operations in all cases.
The gradient of $F$ w.r.t ${\mathbf{T}}$ then reads as
$\nabla_{{\mathbf{T}}}F({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=F_{1}(\overline{{\mathbf{C}}},{\mathbf{T}})+F_{1}(\overline{{\mathbf{C}}}^{\top},{\mathbf{T}})-F_{2}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})-F_{2}({\mathbf{C}}^{\top},\overline{{\mathbf{C}}}^{\top},{\mathbf{T}})$ (76)
When $C_{X}({\mathbf{X}})$ and $C_{Z}({\mathbf{Z}})$ are symmetric, which is the case in all our experiments, this gradient reduces to $\nabla_{{\mathbf{T}}}F=2(F_{1}-F_{2})$.
Low-rank factorization.
Inspired from the work of [65], we propose implementations of $\operatorname{srGW}$ that can leverage the low-rank nature of ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$. Let us assume that both structures can be exactly decomposed as follows, ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{A}}_{1}{\mathbf{A}}_{2}^{\top}$ where ${\mathbf{A}}_{1},{\mathbf{A}}_{2}\in\mathbb{R}^{N\times r}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top}$ with ${\mathbf{B}}_{1},{\mathbf{B}}_{2}\in\mathbb{R}^{n\times s}$, such that $r<<N$ and $s<<n$, can differ respectively from respective dimensions $p$ and $d$ (e.g. for used squared Euclidean distance matrices $r=p+2$ and $s=d+2$ ). For both inner losses $L$ we make use of the following factorization:
$L=L_{2}$: Computing the first term $F_{1}$ coming for the optimized second marginal can benefit from being factored if $d^{2}<<n$. Indeed, as $f_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))={\mathbf{C}}_{Z}({\mathbf{Z}})^{2}=({\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top})\odot({\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top})$, one can use the flattened out product operator described in [65, Section 5], to compute ${\mathbf{C}}_{Z}({\mathbf{Z}})^{2}{\mathbf{T}}^{\top}\bm{1}_{N}={\mathbf{x}}$ in $O(min(n^{2},ns^{2}))$. This way $F_{1}({\mathbf{T}})$ results from stacking $N$ times ${\mathbf{x}}$ in $O(1)$ operations for a total number of computions of $N+O(min(n^{2},ns^{2}))$.And its scalar product with ${\mathbf{T}}$ to compute the loss comes down to $O(Nn)$ additional operations.
Then computing $F_{2}({\mathbf{T}})$ and its scalar product with ${\mathbf{T}}$ can be done following the development of [65, Section 3] for the corresponding GW problem, in $O(Nn(r+s)+rs(N+n))$ operations. So the overall complexity at is $O(Nn(r+s)+rs(N+n)+min(n^{2},ns^{2}))$.
$L=L_{KL}$: In this setting $f_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))={\mathbf{C}}_{Z}({\mathbf{Z}})$ and $h_{1}({\mathbf{C}}_{X}({\mathbf{X}}))={\mathbf{C}}_{X}({\mathbf{X}})$ naturally preserves the low-rank nature of input matrices, but $h_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))=\log({\mathbf{C}}_{Z}({\mathbf{Z}}))$ does not. So computing the first term $F_{1}$, can be performed following this paranthesis order $\bm{1}_{N}((\bm{1}_{N}^{\top}{\mathbf{T}}){\mathbf{A}}_{1})){\mathbf{A}}_{2}^{\top})$ in $O(N(n+s))$ operations. While the second term $F_{2}$ should be computed following this order ${\mathbf{A}}_{1}(({\mathbf{A}}_{2}^{\top}{\mathbf{T}})\log({\mathbf{C}}_{Z}({\mathbf{Z}})))$ in $O(Nnr+rn^{2})$ operations. While their respective scalar product can be computed in $O(Nn)$. So the overall complexity is $O(Nnr+n^{2}r)$.
Notice that in the gaussian kernel case for neighbor embedding methods, where $[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij}=\exp(-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})$ up to some normalization. We have $[h_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))]_{ij}=-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2}$ which admits a low-rank factorization such that we can recover the complexity illustrated above for $L=L_{2}$.
D.2 Solvers.
We develop next our extension of both the Mirror Descent and Conditional Gradient solvers first introduced in [81], for any inner loss $L$ that decomposes as in equation 74 .
Mirror Descent algorithm.
This solver comes down to solve for the exact srGW problem using mirror-descent scheme w.r.t the KL geometry. At each iteration $(i)$, the solver comes down to, first computing the gradient $\nabla_{{\mathbf{T}}}F({\mathbf{T}}^{(i)})$ given in equation 76 evaluated in ${\mathbf{T}}^{(i)}$, then updating the transport plan using the following closed-form solution to a KL projection:
${\mathbf{T}}^{(i+1)}\leftarrow\operatorname{diag}\left(\frac{{\mathbf{h}}}{{\mathbf{K}}^{(i)}\bm{1}_{n}}\right){\mathbf{K}}^{(i)}$ (77)
where ${\mathbf{K}}^{(i)}=\exp\left(\ \nabla_{{\mathbf{T}}}F({\mathbf{T}}^{(i)})-\varepsilon\log({\mathbf{T}}^{(i)})\right)$ and $\varepsilon>0$ is an hyperparameter to tune. Proposition 3 and Lemma 7 in [80, Chapter 6] provides that the Mirror-Descent algorithm converges to a stationary point non-asymptotically when $L=L_{2}$. A quick inspection of the proof suffices to see that this convergence holds for any losses $L$ satisfying equation 74, up to adaptation of constants involved in the Lemma.
Conditional Gradient algorithm.
This algorithm, known to converge to local optimum [39], iterates over the 3 steps summarized in Algorithm 1: .
Algorithm 1 CG solver for $\operatorname{srGW}_{L}$
1:  repeat
2:     ${\mathbf{F}}^{(i)}\leftarrow$ Compute gradient w.r.t ${\mathbf{T}}$ of equation 76.
3:     ${\mathbf{X}}^{(i)}\leftarrow\min_{\begin{subarray}{c}{\mathbf{X}}\bm{1}_{m}={\mathbf{h}}\\ {\mathbf{X}}\geq 0\end{subarray}}\langle{\mathbf{X}},{\mathbf{F}}^{(i)}\rangle$
4:     ${\mathbf{T}}^{(i+1)}\leftarrow(1-\gamma^{\star}){\mathbf{T}}^{(i)}+\gamma^{\star}{\mathbf{X}}^{(i)}$ with $\gamma^{\star}\in[0,1]$ from exact-line search.
5:  until convergence.
This algorithm consists in solving at each iteration $(i)$ a linearization $\langle{\mathbf{X}},{\mathbf{F}}^{(i)}\rangle$ of the problem equation srGW-2 where ${\mathbf{F}}({\mathbf{T}}^{(i)})$ is the gradient of the objective in equation 76. The solution of the linearized problem provides a descent direction ${\mathbf{X}}^{(i)}-{\mathbf{T}}^{(i)}$, and a line-search whose optimal step can be found in closed form to update the current solution ${\mathbf{T}}^{(i)}$. We detail in the following this line-search step for any loss that can be decomposed as in equation 74. It comes down for any ${\mathbf{T}}\in\mathcal{U}_{n}({\mathbf{h}})$, to solve the following problem:
$\gamma=\operatorname*{arg\,min}_{\gamma\in[0,1]}g(\gamma):=F({\mathbf{T}}+\gamma({\mathbf{X}}-{\mathbf{T}}))$ (78)
Observe that this objective function can be developed as a second order polynom $g(\gamma)=a\gamma^{2}+b\gamma+c$. To find an optimal $\gamma$ it suffices to express coefficients $a$ and $b$ to conclude using Algorithm 2 in [78].
Denoting ${\mathbf{X}}^{\top}\bm{1}_{n}={\mathbf{q}}_{X}$ and ${\mathbf{T}}^{\top}\bm{1}_{n}={\mathbf{q}}_{T}$ and following equation 75, we have
$\begin{split}a&=\langle\bm{1}_{n}({\mathbf{q}}_{X}-{\mathbf{q}}_{T})^{\top}f_{2}(\overline{{\mathbf{C}}})^{\top}-h_{1}({\mathbf{C}})({\mathbf{X}}-{\mathbf{T}})h_{2}(\overline{{\mathbf{C}}})^{\top},{\mathbf{X}}-{\mathbf{T}}\rangle\\ &=\langle F_{1}({\mathbf{X}})-F_{1}({\mathbf{T}})-F_{2}({\mathbf{X}})+{\mathbf{F}}({\mathbf{T}}),{\mathbf{X}}-{\mathbf{T}}\rangle\end{split}$ (79)
Finally the coefficient $b$ of the linear term is
$\begin{split}b&=\langle F_{1}({\mathbf{T}})-F_{2}({\mathbf{T}}),{\mathbf{X}}-{\mathbf{T}}\rangle+\langle F_{1}({\mathbf{X}}-{\mathbf{T}})-F_{2}({\mathbf{X}}-{\mathbf{T}}),{\mathbf{T}}\rangle\\ \end{split}$ (80)
Appendix E Appendix of experimental section
We report in the following subsections of this section:
  • E.1: implementation details, validation of hyperparameters, datasets and metrics.
  • E.2: comparison with COOT clustering.
  • E.3: complete scores on all datasets.
  • E.4: study homogeneity vs silhouette score for various numbers of prototypes.
  • E.5: study homogeneity vs k-means score for various numbers of prototypes.
  • E.6: computation time study.
  • E.7: Proofs of concepts with hyperbolic DR kernels.
E.1 Experimental setting
[…]
Implementation. Throughout, the spectral clustering implementation of scikit-learn [56] is used to perform either the clustering steps or the initialization of transport plans. For all methods, ${\mathbf{Z}}$ is initialized from i.i.d. sampling of the standard Gaussian distribution $\mathcal{N}(0,1)$ and further optimized using PyTorch’s automatic differentiation [55] with Adam optimizer [35]. OT-based solvers are built upon the POT [27] library. k-Means is performed using the scikit-learn [56] implementation.
Validated hyperparameters. For the SEA based similarities, we validated perplexity across the set $\{20,50,100,150,200,250\}$. For all kernels, the number of output samples $n$ spans a set of $10$ values, starting at the number of classes in the data and incrementing in steps of $20$. For the computation of ${\mathbf{T}}$ in DistR (see Section 4.1), we benchmark our Conditional Gradient solver, and the Mirror Descent algorithm whose hyperparameter $\varepsilon$ is validated in the two first values within the set $\{10^{i}\}_{i=-3}^{3}$ leading to stable optimization.
Datasets. We provide details about the datasets used in our study. For image datasets, we use COIL-20333https://www1.cs.columbia.edu/CAVE/software/softlib/coil-100.php [53], MNIST and fashion-MNIST444taken from Torchvision [46]. [87]. Regarding single-cell genomics datasets, we rely on PBMC 3k555downloaded from Scanpy [86]. [86], SNAREseq [^6] https://github.com/rsinghlab/SCOT chromatin and gene expression [15] and the scRNA-seq dataset [^7] https://github.com/solevillar/scGeneFit-python from [90] with two hierarchical levels of label. Dimensions are provided in Table 1 When the dimensionality of a dataset exceeds $50$, we pre-process it by applying a PCA in dimension $50$, as done in practice [76].
Table 1: Dataset Sizes.
Number of samples Dimensionality Number of classes
MNIST $10000$ $784$ $10$
F-MNIST $10000$ $784$ $10$
COIL $1440$ $16384$ $20$
SNAREseq (chromatin) $1047$ $19$ $5$
SNAREseq (gene expression) $1047$ $10$ $5$
Zeisel 3005 5000 (8, 49)
PBMC 2638 1838 8
Scores. The homogeneity and NMI scores are taken from Torchmetrics [21]. The other scores used in the experiments are computed as follows.
i) Silhouette: The first step is to associate a label with each prototype ${\mathbf{z}}_{k}$ defined by the $y$ maximizing $\sum_{i\in{[\![N]\!]}}T_{ik}\bm{1}_{y_{i}=y}$ where $y_{i}$ is the class label of the input data point ${\mathbf{x}}_{i}$. Then, to incorporate the relative importance $w_{i}=\left[{\mathbf{h}}_{Z}\right]_{i}$ of each prototype ${\mathbf{z}}_{i}$, we define the weighted mean intra-cluster and nearest-cluster distances that read
$a_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{\sum_{j\in{[\![N]\!]}\delta_{y_{i}=y_{j}}w_{j}d({\mathbf{z}}_{j},{\mathbf{z}}_{j})}}{\sum_{j\in{[\![N]\!]}\delta_{y_{i}=y_{j}}w_{j}}}\quad\text{and}\quad b_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{\sum_{j\in{[\![N]\!]}}\delta_{y_{j}=k}w_{j}d({\mathbf{z}}_{j},{\mathbf{z}}_{i})}{\sum_{j\in{[\![N]\!]}}\delta_{y_{j}=k}w_{j}}$ (81)
such that the silhouette coefficient is $s_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\frac{b_{i}-a_{i}}{\max(b_{i},a_{i})}$ and the final silhouette score reads as
$S({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})=\sum_{i\in{[\![N]\!]}}w_{i}s_{i}({\mathbf{Z}},{\mathbf{Y}},{\mathbf{w}})\>.$ (82)
ii) k-means: We first run the k-means algorithm on the prototypes $\{{\mathbf{z}}_{k}\}_{k\in{[\![n]\!]}}$ giving us the predicted label $\tilde{y}_{k}$ for each $k\in{[\![n]\!]}$. The predicted label for each input data point $i\in{[\![N]\!]}$ is then given by its prototype’s predicted label i.e. $\hat{y}_{i}=\tilde{y}_{\arg\max_{k}T_{ik}}$. Then we compute the NMI score [38] between $(\hat{y}_{i})_{i\in{[\![N]\!]}}$ and the ground-truth class labels $(y_{i})_{i\in{[\![N]\!]}}$.
E.2 Comparison with COOT clustering
The CO-Optimal-Transport (COOT) problem, proposed in [60], is as follows,
$\min_{{\mathbf{T}}_{r}\in{\mathcal{U}}({\mathbf{h}}_{r},\overline{{\mathbf{h}}}_{r})}\min_{{\mathbf{T}}_{c}\in{\mathcal{U}}({\mathbf{h}}_{c},\overline{{\mathbf{h}}}_{c})}\>\sum_{ijkl}(X_{ik}-Z_{jl})^{2}[{\mathbf{T}}_{r}]_{ij}[{\mathbf{T}}_{c}]_{kl}\,,$ (COOT)
where ${\mathbf{h}}_{r}\in\Sigma_{N}$, $\overline{{\mathbf{h}}}_{r}\in\Sigma_{n}$, ${\mathbf{h}}_{c}\in\Sigma_{p}$ and $\overline{{\mathbf{h}}}_{c}\in\Sigma_{d}$. One can seek to optimize the above objective with respect to ${\mathbf{Z}}$ to obtain a competitor method to DistR. This problem is called COOT clustering in [60]. In the latter, ${\mathbf{T}}_{r}$ then plays the role of a soft clustering matrix of the rows of ${\mathbf{X}}$ while ${\mathbf{T}}_{c}$ can be seen as a soft clustering matrix of its columns. The above is thus a linear DR model.
$\mathcal{Z}$ methods ${\mathbf{C}}_{X}$ / ${\mathbf{C}}_{Z}$ COIL MNIST FMNIST PBMC SNA1 SNA2
$\mathbb{R}^{2}$ DistR (ours) SEA / St. 99.50 (0.60) 97.80 (0.00) 93.40 (0.00) 97.10 (0.00) 100.00 (0.00) 100.00 (0.00)
[…]
COOT NA 43.90 (0.70) 9.80 (3.00) 8.50 (0.90) 15.90 (1.90) 44.00 (4.60) 49.70 (8.60)
$\mathbb{R}^{10}$ DistR (ours) $\langle,\rangle_{\mathbb{R}^{p}}$ / $\langle,\rangle_{\mathbb{R}^{d}}$ 96.80 (0.70) 97.00 (0.90) 93.40 (0.00) 97.10 (0.40) 80.50 (0.00) 100.00 (0.00)
[…]
COOT NA 45.50 (1.60) 13.70 (2.10) 9.30 (2.80) 16.10 (2.40) 45.60 (5.90) 76.50 (16.60)
Table 2: Best homogeneity scores for $n$ validated in a span up to $200$ with increments of $20$.
In Table 2, we display the homogeneity values obtained with COOT along with the methods described Section 5. Precisely, it measures to what extent the clustering given by ${\mathbf{T}}_{r}$ groups points with the same ground truth label. One can notice that COOT falls short compared to its competitors that leverage affinity matrices as in state of the art (non-linear) DR methods.
E.3 Complete scores
We complete the results shown in Section 5 by providing the scores obtained on all datasets and models. Scores are plotted in Figure 6 for the t-SNE model and in Figure 7 for the PCA model.
Figure 6: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{2}$) produced by DistR using the t-SNE model: SEA similarity for ${\mathbf{C}}_{X}$ [75], Student’s kernel for ${\mathbf{C}}_{Z}$ and loss $L_{\operatorname{KL}}$.
Figure 7: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{10}$) produced by DistR using the PCA model: $\langle,\rangle_{\mathbb{R}^{p}}$ similarity for ${\mathbf{C}}_{X}$ [75], $\langle,\rangle_{\mathbb{R}^{d}}$ for ${\mathbf{C}}_{Z}$ and loss $L_{2}$.
E.4 Dynamics between homogeneity and silhouette scores across numbers of prototypes
For each method depending on some hyperparameters, models performing the best on average across all numbers of prototypes are illustrated in Figure 8 for spectral methods and in Figure 9 for neighbor embedding ones.
Figure 8: Trade-off between homogeneity vs silhouette scores using PCA model across various numbers of prototypes $n$. The illustration follows the same principal than fig. 3
Figure 9: Trade-off between homogeneity vs silhouette scores using t-SNE model across various numbers of prototypes $n$. The illustration follows the same principal than fig. 3
E.5 Dynamics between homogeneity and NMI scores across numbers of prototypes
For each method depending on some hyperparameters, models performing the best on average across all numbers of prototypes are illustrated in Figure 10 for spectral methods and in Figure 11 for neighbor embedding ones.
Figure 10: Trade-off between homogeneity vs silhouette scores using PCA model across various numbers of prototypes $n$. The illustration follows the same principal than fig. 3
Figure 11: Trade-off between homogeneity vs silhouette scores using t-SNE model across various numbers of prototypes $n$. The illustration follows the same principal than fig. 3
E.6 Computation time comparison
We compare in the following the computation time for all methods benchmarked in Table 1 when using a spectral method and a SNE method. All experiments were done on a server using a GPU (Tesla V100-SXM2-32GB) and composed of 18 cores Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz. All DR steps benefit from our GPU compatible implementation, while spectral clustering (SC) was performed on CPU using scikit-learn implementation running on CPUs.
Notice that to run our experiments we precomputed and saved SC steps for the maximum number of prototypes ran on the input structure $C_{X}({\mathbf{X}})$, for all benchmarked methods e.g CDR used for clustering or DistR and COOT used for initialization of the transport plans. As DRC performs SC over learned embeddings using $C_{Z}({\mathbf{Z}})$ it cannot be precomputed, so we include the time of performing SC for all $n$ for all methods to achieve a fair comparison. Results for three datasets of increasing sample sizes SNA1, ZEISEL and MNIST (See Appendix E.1) are reported in Figure 13 with pre-computation and in Figure 12 without.
Refer to captionRefer to caption
Figure 12: Computation time comparison depending on the number of prototypes $n$ for all methods over 3 datasets, for 5 different initializations, while precomputing spectral clustering on the input structure $C_{X}({\mathbf{X}})$.
Refer to captionRefer to caption
Figure 13: Computation time comparison depending on the number of prototypes $n$ for all methods over 3 datasets, for 5 different initializations, without precomputing spectral clustering on the input structure $C_{X}({\mathbf{X}})$.
Overall our DistR models are competitive with all benchmarked methods in terms of computation time too and can be run (and further validated) in a few seconds using a GPU after precomputing the spectral embeddings for medium size dataset (N=10000 for MNIST), by efficiently leveraging low-rank kernels often used in the DR literature.
E.7 Proofs of concepts with hyperbolic kernels
Hyperbolic spaces [13, 26, 30, 41] are of particular interest as they can capture hierarchical structures more effectively than Euclidean spaces and mitigate the curse of dimensionality by producing representations with lower distortion rates. For instance, [30] adapted t-SNE by using the Poincaré distance and by changing the Student’s t-distribution with a more general hyperbolic Cauchy distribution. Notions of projection subspaces can also be adapted, e.g. [13] use horospheres as one-dimensional subspaces. To match our experiments with the neighbor embeddings in Euclidean settings, we adapt the Symmetric Entropic Affinity (SEA) from [75] for ${\mathbf{C}}_{X}$ and the scalar-normalized student similarity for ${\mathbf{C}}_{Z}$ [76], by simply changing the Euclidean distance by an hyperbolic distance.
Implementation details.
Computations in Hyperbolic spaces are done with Geoopt [36] and the RAdam optimizer [3] replaces Adam. A Wrapped Normal distribution in Hyperbolic spaces [52] is used to initialize ${\mathbf{Z}}$ in the hyperbolic setting. All the computations were conducted in the Lorentz model [54], which is less prone to numerical errors. As such, we used the associated distance function to form ${\mathbf{C}}_{Z}$. After optimization, results are projected back to the Poincaré ball for visualization purpose. In this hyperbolic context, we adopted the formulation of [30] which generalize Student’s t-distributios by Hyperbolic Cauchy distributions (denoted as H-Student in the results). Notice that [30] considered weighted sums of DR objective depending respectively on $L_{2}$ and $L_{KL}$, including 2 additional hyperparamaters, plus various validated curvature levels for the inner hyperbolic distances. In the following experiments we only kept $L_{KL}$ for comparison with the Euclidean SNE based methods illustrated in Section 5 and previous Sections of E, while validating the same hyperparameters and setting the space curvature to $1$. The silhouette score was adapted to this kernel considering the Hyperbolic distance insteand of the euclidean one, and we implemented a Hyperbolic Kmeans whose barycenters are estimated using the RAdam optimizer to compute the NMI scores.
Results.
[…]
Refer to caption Refer to caption
Figure 14: Best trade-off between homogeneity vs silhouette (2 first plots), and homogeneity vs NMI (2 last plots). Scores are normalized in $\left[0,1\right]$ via min-max scaling over a dataset. Small markers represent scores for 5 runs for a given dataset, while big ones are their mean. For each method we illustrate the 20-80% percentiles of normalized scores as a colored surface.
[…]
Figure 15: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{10}$) produced by DistR using the Hyperbolic Student model.